### v3

# Combining exchangeable p-values

Abstract
The problem of combining p-values is an old and fundamental one, and the classic assumption of independence is often violated or unverifiable in many applications. There are many well-known rules that can combine a set of arbitrarily dependent p-values (for the same hypothesis) into a single p-value. We show that essentially all these existing rules can be strictly improved when the p-values are exchangeable, or when external randomization is allowed (or both). For example, we derive randomized and/or exchangeable improvements of well known rules like “twice the median” and “twice the average”, as well as geometric and harmonic means. Exchangeable p-values are often produced one at a time (for example, under repeated tests involving data splitting), and our rules can combine them sequentially as they are produced, stopping when the combined p-values stabilize. Our work also improves rules for combining arbitrarily dependent p-values, since the latter becomes exchangeable if they are presented to the analyst in a random order. The main technical advance is to show that all existing combination rules can be obtained by calibrating the p-values to e-values (using an $\alpha$-dependent calibrator), averaging those e-values, converting to a level-$\alpha$ test using Markov’s inequality, and finally obtaining p-values by combining this family of tests; the improvements are delivered via recent randomized and exchangeable variants of Markov’s inequality.
1 Introduction
The combination of p-values represents a fundamental task frequently encountered in statistical inference and its applications in the natural sciences. Within the realm of multiple testing, for instance, the focus lies in testing whether all individual null hypotheses are simultaneously true. This particular challenge, often referred to as global null testing, can be addressed by merging multiple p-values into a single p-value. Potential solutions, based on the independence of p-values, are provided in Fisher, (1925), Pearson, (1934), and Simes, (1986), with the latter also working under a certain notion of positive dependence (Sarkar,, 1998; Benjamini and Yekutieli,, 2001). See Owen, (2009) for a review of the methods. Recently, harmonic mean p-values (Wilson,, 2019) and methods under negative dependence (Chi et al.,, 2024) have been developed in the context of multiple testing.
The assumption of independence (or positive/negative dependence) is often violated in many real-world applications (see e.g., Section 4.2 of Efron,, 2010), and in certain scenarios, it may be preferable not to impose unverifiable conditions on the joint distribution of the p-values, beyond the minimum necessary assumption that each individual p-value is indeed (marginally) valid. Several methods are available for combining p-values with arbitrary dependencies; notably, the Bonferroni method is widely used, which involves multiplying the minimum of the p-values by the number of tests conducted. Other methods have been proposed in the literature, some based on order statistics (Rüger,, 1978; Morgenstern,, 1980; Hommel,, 1983), while others rely on their arithmetic mean and other variants (Rüschendorf,, 1982; Vovk and Wang,, 2020; Vovk et al.,, 2022). Two prominent examples are that both 2 times the median of the p-values, and 2 times the average of the p-values, are valid combination rules, and the multiplicative factor of 2 cannot be reduced. Inevitably, these methods that assume arbitrary dependence come with a price to pay in terms of statistical power. Our work will improve all of these rules under the weaker assumption of exchangeability.
To elaborate, the main objective of our work is to obtain new valid merging methods under the assumption of exchangeability of the input p-values, that are more powerful than methods that assume arbitrary dependence. Implied by the relative strength of the dependence assumptions, the new methods will be incomparable to methods assuming negative or positive dependence, and less powerful than methods assuming independence. However, specialized methods for handling exchangeable dependence are quite practically relevant. Such dependence is encountered, for example, in statistical testing via sample splitting. There are at least two different reasons for which sample splitting is used: the first is to relax the assumptions needed to obtain theoretical guarantees, while the second is to reduce computational costs. Some examples of such procedures are Cox, (1975); Wasserman and Roeder, (2009); Banerjee et al., (2019); Wasserman et al., (2020). The drawback of these methods based on sample-splitting is that the obtained p-values are affected by the randomness of the split. Meinshausen et al., (2009) called this phenomenon as a p-value lottery. The problem of merging p-values from different data splits was studied in DiCiccio et al., (2020), but our combination rules are more powerful, and also more general and systematic because they apply more broadly. The same problem has also been studied in Guo and Shah, (2023), where the authors propose a method based on subsampling to merge test statistics based on different random splits.
Choi and Kim, (2023) showed that in the aforementioned rule of “twice the average”, the constant factor of 2 cannot be improved even under exchangeability. Importantly, their result does not imply that “twice the average” cannot be improved; it simply states that any such improvement cannot proceed by attempting to lower the constant of 2. Indeed, our paper will improve on this well known rule (and many others), but it proceeds differently.
It is perhaps interesting that all the aforementioned methods for merging under arbitrary dependence or under exchangeability are actually inadmissible when randomization is permitted. Randomization in the context of hypothesis testing is not new and is used, for example, in discrete tests; some examples are Fisher’s exact test (Fisher,, 1925) or the randomized test for a binomial proportion proposed in Stevens, (1950). In our paper, we will see how the introduction of a simple external randomization (an independent uniform random variable and/or uniform permutation) can improve the existing merging rules for arbitrary dependence, as well as our new rules for exchangeable merging.
In terms of technical aspects, one of our main contributions is to point out explicitly how existing merging rules for arbitrary dependence are actually recovered in a unified manner: by transforming the p-values into e-values (Vovk and Wang,, 2021; Wasserman et al.,, 2020; Grünwald et al.,, 2024) using different “calibrators”, averaging the resulting e-values and finally applying Markov’s inequality. This connection is particularly important, because then the improvements under exchangeability, or that provided by randomization, are achieved by invoking the recent “exchangeable Markov’s inequality” and “uniformly randomized Markov’s inequality” (Ramdas and Manole,, 2023).
Paper outline and peek at results.
The rest of this paper is organized as follows. In Section 2, we introduce the notation and tools necessary for the paper. In Section 3, the main results are presented in a general way, focusing on two distinct aspects: the first part addresses the case of exchangeable p-values, while the second part introduces novel findings under the assumption of arbitrarily dependent p-values when randomization is allowed. Subsequent sections investigate the implications of these results across various p-merging functions commonly found in the literature. Specifically, Section 4 and Section 5 delve into the combination proposed by Rüger, (1978) and Hommel, (1983), respectively. Section 6 examines the case of arithmetic mean. The following two sections address two additional scenarios within the family of generalized means: namely, the harmonic mean (Section 7) and the geometric mean (Section 8). Section 9 presents some simulation results, before we conclude in Section 10.
Before proceeding with the paper, Table 1 presents some notable combination rules introduced in the literature and their corresponding exchangeable and randomized versions introduced in the following sections. These results are first derived in a general form and then discussed case-by-case. Some of the rules in the table are not admissible, as will be explained in the following.
Combination rule Arbitrary dependence (known) Exchangeability (new) Arbitrary dependence, randomized (new)
Rüger combination $\frac{K}{k}p_{(k)}$ $\frac{K}{k}\bigwedge_{m=1}^{K}p^{m}_{(\lambda_{m})}$ $\frac{K}{k}p_{(\lceil uk\rceil)}$
Arithmetic mean $2A(\mathbf{p})$ $2\bigwedge_{m=1}^{K}A(\mathbf{p}_{m})$ $\frac{2}{2-u}A(\mathbf{p})$
Geometric mean $eG(\mathbf{p})$ $e\bigwedge_{m=1}^{K}G(\mathbf{p}_{m})$ $e^{u}G(\mathbf{p})$
Harmonic mean $(T_{K}+1)H(\mathbf{p})$ $(T_{K}+1)\bigwedge_{m=1}^{K}H(\mathbf{p}_{m})$ $(T_{K}u+1)H(\mathbf{p})$
Table 1: Some combination rules for arbitrarily dependent p-values documented in literature, along with their exchangeable and randomized improvements. If randomization is permitted, one can also improve the existing rules for combining arbitrarily dependent p-values by using the exchangeable combination rule applied to a random permutation of the p-values (this is not presented as a separate column). Here, $\mathbf{p}=(p_{1},\dots,p_{K})$ denotes the vector of p-values, and $\mathbf{p}_{m}$ represents the vector containing the first $m$ values of $\mathbf{p}$. In the table, $p_{(k)}$ is the $k$-th smallest value of $\mathbf{p}$, while $p_{(\lambda_{m})}^{m}$ is the $\lambda_{m}=\lceil m\frac{k}{K}\rceil$ ordered value of $\mathbf{p}_{m}$. The value $u$ is the realization of a uniform random variable on $[0,1]$. Additionally, $A,G$ and $H$ respectively denote the arithmetic mean, the geometric mean, and the harmonic mean. The value $T_{K}$ is given by $T_{K}=\log K+\log\log K+1$, $K\geq 2$.
2 Problem setup and notation
Without loss of generality, let $(\Omega,\mathcal{F},\mathbb{P})$ be an atomless probability space [^1] A probability space is atomless if there exists a uniform random variable on this space, and this is implicitly assumed in almost all papers in statistics; see Vovk and Wang, (2021, Appendix D) for related results and discussions. and $\mathcal{U}$ be the set of all uniform random variables on $[0,1]$ under $\mathbb{P}$. In the following, $K\geq 2$ is an integer.
A p-variable for testing $\mathbb{P}$ is a random variable $P:\Omega\to[0,\infty)$ satisfying
$\mathbb{P}(P\leq\alpha)\leq\alpha,$
for all $\alpha\in(0,1)$. Typically, of course, $P$ will only take values in $[0,1]$, but nothing is lost by allowing the larger range above. For all results on validity of the methods in this paper, it suffices to consider p-variables in $\mathcal{U}$, i.e., $\mathbb{P}(P\leq\alpha)=\alpha$ for each $\alpha\in(0,1)$.
An e-variable for testing $\mathbb{P}$ is a non-negative extended random variable $E:\Omega\to[0,\infty]$ with $\mathbb{E}_{\mathbb{P}}[E]\leq 1$. A calibrator is a decreasing function $f:[0,\infty)\to[0,\infty]$ satisfying $f=0$ on $(1,\infty)$ and $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$. Essentially, a calibrator transforms any p-variable to an e-variable. It is admissible if it is upper semicontinuous, $f(0)=\infty$, and $\int_{0}^{1}f(p)\mathrm{d}p=1$. Equivalently, a calibrator is admissible if it is not strictly dominated, in a natural sense, by any other calibrator (Proposition 2.1 and Proposition 2.2 in Vovk and Wang,, 2021). We fix $\mathbb{P}$ throughout, and omit “for testing $\mathbb{P}$” when discussing p-variables and e-variables; we do not distinguish them from the commonly used terms “p-values” and “e-values”, and this should create no confusion.
Our starting point is a collection of $K$ p-variables $\mathbf{P}=(P_{1},\dots,P_{K})$ and we denote their observed (realized) values by $\mathbf{p}=(p_{1},\dots,p_{K})$. Borrowing terminology from Vovk et al., (2022) and Vovk and Wang, (2020), we have that a p-merging function is an increasing Borel function $F:[0,\infty)^{K+1}\to[0,\infty)$ such that $\mathbb{P}(F(\mathbf{P})\leq\alpha)\leq\alpha$ whenever $P_{1},\dots,P_{K}$ are p-variables. In other words, the function $F$, starting from $K$ p-values, returns a valid p-value. A p-merging function is symmetric if $F(\mathbf{p})$ is invariant under any permutation of $\mathbf{p}$, and it is homogeneous if $F(\gamma\mathbf{p})=\gamma F(\mathbf{p})$ for all $\mathbf{p}$ with $F(\mathbf{p})\leq 1$ and $\gamma\in(0,1]$. The class of homogeneous p-merging functions encompasses the O-family based on quantiles introduced in Rüger, (1978), the Hommel’s combination and the M-family introduced in Vovk and Wang, (2020). We now introduce the notion of domination in the context of p-merging functions.
Definition 2.1.
A function $F$ dominates (interpreted as better being smaller) a function $G$ if
$F(\mathbf{p})\leq G(\mathbf{p}),\quad\mathrm{for\leavevmode\nobreak\ all\leavevmode\nobreak\ }\mathbf{p},$ (1)
and the domination is strict if $F(\mathbf{p})<G(\mathbf{p})$, for at least one $\mathbf{p}$. A p-merging function $F$ is admissible if it is not strictly dominated by any other p-merging function.
Although we defined p-values and e-values for a single probability measure $\mathbb{P}$, all results hold for composite hypotheses. More precisely, if $\mathbf{P}$ is a vector of p-variables for a composite hypothesis and $F$ is a p-merging function, then $F(\mathbf{P})$ is a p-variable for the same composite hypothesis. See Vovk and Wang, (2021) for precise definitions and related discussions.
For any function $F:[0,\infty)^{K}\to[0,\infty)$ and $\alpha\in(0,1)$, let its rejection region at level $\alpha$ be given by
$R_{\alpha}(F):=\left\{\mathbf{p}\in[0,\infty)^{K}:F(\mathbf{p})\leq\alpha\right\}.$ (2)
For any homogeneous $F$ and $\alpha\in(0,1)$, $R_{\alpha}(F)$ takes the form $R_{\alpha}(F)=\alpha A$ for some $A\subseteq[0,\infty)^{K}$. Conversely, any increasing collection of Borel lower sets $\{R_{\alpha}\subseteq[0,\infty)^{K}:\alpha\in(0,1)\}$ determines an increasing Borel function $F:[0,\infty)^{K}\to[0,1]$ by the equation
$F(\mathbf{p})=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\},$ (3)
with the convention $\inf\varnothing=1$. It is immediate to see that $F$ is a p-merging function if and only if $\mathbb{P}(\mathbf{P}\in R_{\alpha})\leq\alpha$ for all $\alpha\in(0,1)$ and $\mathbf{P}\in\mathcal{U}^{K}$.
Below, $\Delta_{K}$ is the standard $K$-simplex. Every admissible homogeneous p-merging function possesses a dual formulation expressed in terms of calibrators, as summarized below.
Theorem 2.2 (Vovk et al., (2022); Theorem 5.1).
For any admissible homogeneous p-merging function $F$, there exist $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$ and admissible calibrators $f_{1},\dots,f_{K}$ such that
$R_{\alpha}(F)=\alpha\left\{\mathbf{p}\in[0,\infty)^{K}:\sum_{k=1}^{K}\lambda_{k}f_{k}(p_{k})\geq 1\right\}\qquad\mathrm{for\leavevmode\nobreak\ each\leavevmode\nobreak\ \alpha\in(0,1)}.$ (4)
Conversely, for any $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$ and calibrators $f_{1},\dots,f_{K}$, (4) determines a homogeneous p-merging function.
We will exploit this dual form to implement our “randomized” and “exchangeable” techniques, generating p-merging functions that consistently give smaller p-values (usually strictly) than those produced by their original counterparts. From (4), it is worth noting that $\sum_{k=1}^{K}\lambda_{k}f_{k}(P_{k})$ is a valid e-value. We now define a lemma that will be useful in the following.
Lemma 2.3.
Let $f_{1},\dots,f_{K}$ be $K$ calibrators and $\mathbf{P}\in\mathcal{U}^{K}$. Then, for any $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$ and $\alpha\in(0,1]$, we have that
$\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)$ (5)
is an e-value. If $f_{1},\dots,f_{K}$ are admissible calibrators, then $\mathbb{E}\left[\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\right]=1.$
Proof.
By definition, the quantity in (5) is non-negative. In addition,
$\begin{split}\mathbb{E}\left[\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\right]&=\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}\mathbb{E}\left[f_{k}\left(\frac{P_{k}}{\alpha}\right)\right]=\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}\int_{0}^{\alpha}f_{k}\left(\frac{p}{\alpha}\right)\mathrm{d}p\\ &=\sum_{k=1}^{K}\lambda_{k}\int_{0}^{1}f_{k}\left(p\right)\mathrm{d}p\leq 1.\end{split}$
If the calibrators are admissible, one can see that the equality holds since $\int_{0}^{1}f_{k}\left(p\right)\mathrm{d}p=1$ for each $k$. ∎
In particular, choosing $\lambda_{1}=1$ and $\lambda_{k}=0$, for $k\geq 2$, we have that $(1/\alpha)f_{1}(P_{1}/\alpha)$ is an e-value, for all $\alpha\in(0,1]$.
Before introducing our results, we present some inequalities introduced in Ramdas and Manole, (2023) that will be fundamental throughout the subsequent discussion. The following inequalities can be viewed as an extension of the Markov inequality.
Theorem 2.4 (Exchangeable Markov Inequality).
Let $X_{1},X_{2},\dots$ form an exchangeable sequence of non-negative and integrable random variables. Then, for any $a>0$,
$\mathbb{P}\left(\exists k\geq 1:\frac{1}{k}\sum_{i=1}^{k}X_{i}\geq\frac{1}{a}\right)\leq a\mathbb{E}[X_{1}].$
The second inequality is based on an external randomization of the threshold of the Markov’s inequality.
Theorem 2.5 (Uniformly-randomized Markov Inequality).
Let $X$ be a non-negative random variable independent of $U\in\mathcal{U}$. Then, for any $a>0$,
$\mathbb{P}(X\geq U/a)\leq a\mathbb{E}[X].$
The third inequality combines the previous two theorems in the following way:
Theorem 2.6 (Exchangeable and uniformly-randomized Markov Inequality).
Let $X_{1},\dots,X_{K}$ be a set of exchangeable and non-negative random variables independent of $U\in\mathcal{U}$. Then, for any $a>0$,
$\mathbb{P}\left(X_{1}\geq U/a\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}X_{i}\geq\frac{1}{a}\right)\leq a\,\mathbb{E}[X_{1}].$
These results will be used in the next section as technical tools to derive new combination rules in different situations.
3 New results on merging p-values
This section introduces our main results stated in abstract and general terms, which we instantiate in special cases (like arithmetic or geometric mean) in the sections that follow.
3.1 Exchangeable p-values
Assuming i.i.d. p-values is often overly stringent in numerous practical applications. A more pragmatic and less restrictive assumption is exchangeability of p-values, indicating that the distribution of the p-values is unchanged under a random permutation of the indices. Formally,
$(P_{1},\dots,P_{K})\stackrel{{\scriptstyle d}}{{=}}(P_{\sigma(1)},\dots,P_{\sigma(K)}),$
where =$\stackrel{{\scriptstyle d}}{{=}}$ represents equality in distribution while $\sigma:\{1,\dots,K\}\to\{1,\dots,K\}$ is any permutation of the indices. As discussed in Section 1, this situation is encountered in statistical testing using repeated sample splitting (repeated $K$ times on the same data in an identical fashion). In this section, we assume that the sequence of p-variables $\mathbf{P}=(P_{1},\dots,P_{K})$ is exchangeable and takes values in $[0,1]^{K}$.
Remark 3.1.
The reader may note that exchangeability can be induced by processing the (potentially non-exchangeable) sequence of p-values $(P_{1},\dots,P_{K})$ in a uniformly random order. As a consequence, this implies that if randomization is allowed, it is always possible to satisfy the exchangeability assumption even if the starting sequence has an arbitrary dependence. Stated alternatively, exchangeable combination rules can be safely applied to arbitrarily dependent p-values if the p-values are presented to the analyst in a random order.
We now present an extension of Theorem 2.2 under exchangeability.
Theorem 3.2.
Let $f$ be a calibrator, and $\mathbf{P}=(P_{1},\dots,P_{K})\in\mathcal{U}^{K}$ be exchangeable. For each $\alpha\in(0,1)$, we have
$\mathbb{P}\left(\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\alpha}\right)\geq 1\right)\leq\alpha.$
Proof.
The proof involves the use of the exchangeable Markov inequality (EMI) recalled in Theorem 2.4:
$\mathbb{P}\left(\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\alpha}\right)\geq 1\right)\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\left[f\left(\frac{P_{1}}{\alpha}\right)\right]=\alpha\mathbb{E}\left[\frac{1}{\alpha}f\left(\frac{P_{1}}{\alpha}\right)\right]\stackrel{{\scriptstyle(ii)}}{{\leq}}\alpha,$
where $(i)$ is due to EMI while $(ii)$ holds due to Lemma 2.3. ∎
The result obtained in the previous theorem gives the intuition to derive better p-merging functions by exploiting the duality between rejection regions and p-merging functions. First, we define an exchangeable p-merging function.
Definition 3.3.
An exchangeable p-merging function is an increasing Borel function $F:[0,1]^{K}\to[0,1]$ such that $\mathbb{P}(F(\mathbf{P})\leq\alpha)\leq\alpha$ for all $\alpha\in(0,1)$ and $\mathbf{P}\in\mathcal{U}^{K}$ that is exchangeable. It is homogeneous if $F(\gamma\mathbf{p})=\gamma F(\mathbf{p})$ for all $\gamma\in(0,1]$ and $\mathbf{p}\in[0,1]^{K}$. An exchangeable p-merging function $F$ is admissible if for any exchangeable p-merging function $G$, $G\leq F$ implies $G=F$.
Starting from a calibrator $f$ and $\alpha\in(0,1)$, we define the exchangeable rejection region by
$R_{\alpha}=\left\{\mathbf{p}\in[0,1]^{K}:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\mathrm{\leavevmode\nobreak\ holds\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }k\leq K\right\}.$
Using $R_{\alpha}$, we can define the function $F:[0,1]^{K}\to[0,1]$ by
$\begin{split}F(\mathbf{p})&=\inf\{\alpha\in(0,1):\mathbf{p}\in R_{\alpha}\}\\ &=\inf\left\{\alpha\in(0,1):\exists k\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\right\}.\\ &=\inf\left\{\alpha\in(0,1):\bigvee_{k=1}^{K}\left(\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\right)\geq 1\right\},\end{split}$ (6)
with the convention $\inf\varnothing=1$. Note that (6) is always smaller or equal than the combination
$F^{\prime}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}f\left(\frac{p_{k}}{\alpha}\right)\geq 1\right\},$
which is valid for p-values with an arbitrary dependence. Indeed, all admissible homogeneous and symmetric p-merging functions have the form $F^{\prime}$ for some admissible calibrator (Vovk et al.,, 2022, Theorem 5.2).
Theorem 3.4.
If $f$ is a calibrator and $\mathbf{P}\in\mathcal{U}^{K}$ is an exchangeable sequence, then $F$ in (6) is an exchangeable homogeneous p-merging function.
Proof.
It is clear that $F$ is increasing and Borel since $R_{\alpha}$ is a lower set. For an exchangeable sequence $\mathbf{P}\in\mathcal{U}^{K}$ and $\alpha\in(0,1)$, using Theorem 3.2 and the fact that $(R_{\beta})_{\beta\in(0,1)}$ is nested, we have
$\begin{split}\mathbb{P}\left(F(\mathbf{P})\leq\alpha\right)&=\mathbb{P}\big{(}\inf\{\beta\in(0,1):\mathbf{P}\in R_{\beta}\}\leq\alpha\big{)}\\ &=\mathbb{P}\left(\inf\left\{\beta\in(0,1):\exists k\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\geq 1\right\}\leq\alpha\right)\\ &=\mathbb{P}\left(\bigcap_{\beta>\alpha}\left\{\exists k\leq K:\left(\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\right)\geq 1\right\}\right)\\ &=\inf_{\beta>\alpha}\mathbb{P}\left(\exists k\leq K:\left(\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\right)\geq 1\right)\leq\inf_{\beta>\alpha}\beta=\alpha.\end{split}$
Therefore $F$ is a valid p-merging function. Homogeneity comes directly from the definition of (6). ∎
It is clear from (6) that the function depends on the order of the values in $\mathbf{p}$, and hence $F(\mathbf{p})$ is not a symmetric function. This is not a coincidence: in the next result, we show that any symmetric exchangeable p-merging function is actually valid under arbitrary dependence, and hence they cannot improve over the admissible p-merging functions studied by Vovk et al., (2022). In particular, this implies that under exchangeability the multiplicative factor 2 for the arithmetic average cannot be improved, as earlier noted by Choi and Kim, (2023), but it also extends their result to every other symmetric merging function.
Proposition 3.5.
A symmetric exchangeable p-merging function is necessarily a p-merging function. Hence, for an exchangeable p-merging function to strictly dominate an admissible p-merging function, it cannot be symmetric.
Proof.
Let $\mathbf{P}\in\mathcal{U}^{K}$, and let $\sigma$ be a random permutation of $\{1,\dots,K\}$, uniformly drawn from all permutations of $\{1,\dots,K\}$ and independent of $\mathbf{P}$. Let $\mathbf{P}^{\sigma}=(P_{\sigma(1)},\dots,P_{\sigma(K)})$. Note that $\mathbf{P}^{\sigma}$ is exchangeable by construction. If $F$ is a symmetric exchangeable p-merging function, it must satisfy $F(\mathbf{P}^{\sigma})=F(\mathbf{P})$. Because $F(\mathbf{P}^{\sigma})$ is a p-variable, so is $F(\mathbf{P})$, showing that $F$ is a p-merging function. ∎
Clearly, Proposition 3.5 implies that under symmetry, a function is a p-merging function if and only if it is an exchangeable p-merging function. Hence, to take advantage of the exchangeability of the p-values (over arbitrary dependence), one necessarily deviates from symmetric ways of merging p-values, as done in Theorem 3.4. More importantly, the proof of Proposition 3.5 illustrates the idea (mentioned in Remark 3.1) that for arbitrarily dependent p-values, we can first randomly permute them and then apply an exchangeable p-merging function (not necessarily symmetric) such as the one in Theorem 3.4, to obtain a valid p-value.
The next result gives a simple condition on the calibrator $f$ that guarantees that the probability of rejection using $F$ in (6) is sharp for some $\mathbf{P}$.
Proposition 3.6.
Suppose that $f$ is a convex admissible calibrator with $f(0+)\leq K$ and $f(1)=0$, and $F$ is in (6). For $\alpha\in(0,1)$, there exists an exchangeable $\mathbf{P}\in\mathcal{U}^{K}$ such that $\mathbb{P}(F(\mathbf{P})\leq\alpha)=\alpha.$
Proof.
Take $U\in\mathcal{U}$ and an event $A$ with $\mathbb{P}(A)=\alpha$ independent of $U$. Let $b=f(0+)\leq K$. The condition on $f$ guarantees that $f(U)$ is a random variable with support $[0,b]$, mean $1$ and a decreasing density. The above conditions, using Theorem 3.2 of Wang and Wang, (2016), guarantee that there exists $\mathbf{U}=(U_{1},\dots,U_{K})\in\mathcal{U}^{K}$ such that $\mathbb{P}(\sum_{i=1}^{K}f(U_{i})=K)=1$. We assume that $U,A,\mathbf{U}$ are mutually independent; this is possible as we are only concerned with distributions. Taking a uniformly drawn random permutation further allows us to assume that $\mathbf{U}$ is exchangeable. Let $P_{i}=\alpha U_{i}\mathds{1}_{A}+(\alpha+(1-\alpha)U)\mathds{1}_{A^{c}}$ for $i=1,\dots,K$. It is clear that each $P_{i}\in\mathcal{U}$ and $\mathbf{P}=(P_{1},\dots,P_{K})$ is exchangeable. Moreover, by the definition of $F$,
$\mathbb{P}(F(\mathbf{P})\leq\alpha)\geq\mathbb{P}\left(\frac{1}{K}\sum_{i=1}^{K}f(P_{i}/\alpha)\geq 1\right)=\mathbb{P}(A)\mathbb{P}\left(\sum_{i=1}^{K}f(U_{i})=K\right)=\alpha.$
The other inequality $\mathbb{P}(F(\mathbf{P})\leq\alpha)\leq\alpha$ follows from Theorem 3.2. ∎
Remark 3.7.
Proposition 3.6 is not sufficient to justify admissibility of $F$ in (6). In general, admissibility of exchangeable p-merging functions remains unclear. For instance, take $F$ in (6) with $f(p)=(2-2p)_{+}$, corresponding to the arithmetic average, as in Section 6 below. If $K=2$, then $F$ is not admissible as it is strictly dominated by the Bonferroni p-merging function given by $F_{\mathrm{Bonf}}(p_{1},\dots,p_{K})=K\min\{p_{1},\dots,p_{K}\}$. For $K\geq 3$, $F$ and $F_{\mathrm{Bonf}}$ are not comparable.
3.2 Sequentially combining a stream of exchangeable p-values
In Section 3.1 we have seen that, if our starting vector of p-values $\mathbf{P}=(P_{1},\dots,P_{K})$ is exchangeable, then it is possible to derive new combination rules exploiting the exchangeability of the p-values. The technique for developing these new rules involves converting the initial p-values into e-values through an $\alpha$-dependent calibrator. Following this transformation, the exchangeable Markov inequality (Theorem 2.4) plays a crucial role in the formulation of these rules. In particular, we can notice that the inequality in Theorem 2.4 is uniformly valid; therefore, the exchangeable sequence of p-values need not be limited to a set with cardinality $K$, but it is possible to continue to add p-values and stop when the procedure is stable.
To provide a concrete example, in the case where p-values are obtained by applying a sample-splitting procedure to the same dataset, a researcher can obtain one p-value at a time simply by performing a new data split. The researcher would then aim to combine these p-values sequentially and potentially stop when the procedure appears to stabilize.
We first define the following lemma:
Lemma 3.8.
Let $\mathbf{p}=(\mathbf{p}_{1},\mathbf{p}_{2})\in[0,1]^{K}$ be a vector with $\mathbf{p}_{1}\in[0,1]^{K_{1}},\mathbf{p}_{2}\in[0,1]^{K_{2}}$ and $K=K_{1}+K_{2}$. In addition, let $F$ be the function defined in (6). Then
$F(\mathbf{p})=F(\mathbf{p}_{1},\mathbf{p}_{2})\leq F(\mathbf{p}_{1}).$
Proof.
Fix any $\mathbf{p}=(\mathbf{p}_{1},\mathbf{p}_{2})\in[0,1]^{K}$ and suppose that $F(\mathbf{p}_{1})=\alpha$. By definition, we have
$\exists k\leq K_{1}:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\implies\exists k\leq K_{1}+K_{2}:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1.$
This implies
$F(\mathbf{p})=\inf\left\{\beta\in(0,1):\exists k\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\beta}\right)\geq 1\right\}\leq\alpha,$
due to the fact that $f$ is increasing in $\beta$. ∎
The above result implies that the function $F$ in (6) is non-increasing as more parameters (i.e., p-values) are added. In addition, the following holds.
Theorem 3.9.
Let $P_{1},P_{2},...\in\mathcal{U}^{\infty}$ be an infinite exchangeable sequence and let $F$ be the function defined in (6). Then,
$\mathbb{P}(\exists k\geq 1:F(\mathbf{P}_{k})\leq\alpha)\leq\alpha,$
where $\mathbf{P}_{k}=(P_{1},\dots,P_{k})$.
Proof.
From Lemma 3.8, we have that
$\exists k\leq K:F(\mathbf{p}_{k})\leq\alpha\iff F(\mathbf{p}_{K})\leq\alpha,$
where $\mathbf{p}_{k}=(p_{1},\dots,p_{k})$ is the sequence continaing the first $k$ values of $\mathbf{p}$. Then we can write,
$\begin{split}\mathbb{P}\left(\exists k\geq 1:F(\mathbf{P}_{k})\leq\alpha\right)&=\mathbb{P}\left(\bigcup_{K\in\mathbb{N}}\{\exists k\leq K:F(\mathbf{P}_{k})\leq\alpha\}\right)\\ &=\lim_{K\to\infty}\mathbb{P}\left(\exists k\leq K:F(\mathbf{P}_{k})\leq\alpha\right)\\ &=\lim_{K\to\infty}\mathbb{P}\left(F(\mathbf{P}_{K})\leq\alpha\right)\leq\alpha,\end{split}$
where the last inequality is due to Theorem 3.4. ∎
The result allows the analyst to either continue collecting new (exchangeable) p-values or to cease based on the outcome. In particular, in the example described at the beginning of the section, a researcher can stop the procedure when the result seems to stabilize. However, there are some issues to consider when applying such a procedure. In particular, some calibrators $f$ depend on the number $K$ of p-values. For example, this is the case for the Hommel combination (Section 5) and for the harmonic mean (Section 7). As a final caveat, exchangeability becomes more stringent when the number $K$ grows; for instance, in general, for a given $K$-dimensional exchangeable vector $(P_{1},\dots,P_{K})$, there may not exist $P_{K+1}$ such that $(P_{1},\dots,P_{K+1})$ is exchangeable. Luckily, in many practical situations, the procedure that produced the $K$ exchangeable p-values in the first place could also produce more of them; for example, this happens when the p-value was produced by sample splitting.
3.3 Randomized p-merging functions
In this subsection, we start with a collection of arbitrarily dependent p-values and we will show how it is possible to enhance existing merging rules using a simple randomization trick. In this case, we denote by
$\mathcal{U}^{K}\otimes\mathcal{U}=\{(\mathbf{P},U)\in\mathcal{U}^{K}\times\mathcal{U}:\mbox{$U$ and $\mathbf{P}$ are independent}\},$
and we state a randomized version of the converse direction of Theorem 2.2, by changing the constant $1$ in (4) to a uniform random variable $U$.
Theorem 3.10.
Let $f_{1},\dots,f_{K}$ be calibrators and $(P_{1},\dots,P_{K},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$. For each $\alpha\in(0,1)$ and $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$, we have
$\mathbb{P}\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\geq U\right)\leq\alpha.$ (7)
If $f_{1},\dots,f_{K}$ are admissible calibrators and $\mathbb{P}(\sum_{k=1}^{K}\lambda_{k}f_{k}(P_{k}/\alpha)\leq 1)=1$, then equality holds
$\mathbb{P}\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\beta}\right)\geq U\right)=\beta\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ for all $\beta\in(0,\alpha]$}.$ (8)
Proof.
From direct calculation and using Lemma 2.3,
$\mathbb{P}\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\geq U\right)$ $=\mathbb{E}\left[\mathbb{P}\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\geq U\right)\mid\mathbf{P}\right]$
$=\mathbb{E}\left[\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\right)\wedge 1\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\right]=\alpha\mathbb{E}\left[\frac{1}{\alpha}\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\alpha}\right)\right]\leq\alpha,$
The equality for $\beta=\alpha$ follows because $\sum_{k=1}^{K}\lambda_{k}f_{k}(P_{k}/\alpha)\leq 1$ and $\int_{0}^{1}f_{k}(p)\mathrm{d}p=1$ for each $k$ guarantee the inequalities in the above set of equations are equalities. For $\beta<\alpha$, it suffices to notice that $\sum_{k=1}^{K}\lambda_{k}f_{k}(P_{k}/\alpha)$ is increasing in $\alpha$. ∎
The result in Theorem 3.10 is a direct consequence of the uniformly randomized Markov inequality (UMI) introduced by Ramdas and Manole, (2023); see Theorem 2.5.
Definition 3.11.
A randomized p-merging function is an increasing Borel function $F:[0,1]^{K+1}\to[0,1]$ such that $\mathbb{P}(F(\mathbf{P},U)\leq\alpha)\leq\alpha$ for all $\alpha\in(0,1)$ and $(\mathbf{P},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$. It is homogeneous if $F(\gamma\mathbf{p},u)=\gamma F(\mathbf{p},u)$ for all $\gamma\in(0,1]$ and $(\mathbf{p},u)\in[0,1]^{K+1}$. A randomized p-merging function $F$ is admissible if for any randomized p-merging function $G$, $G\leq F$ implies $G=F$.
Let $f_{1},\dots,f_{K}$ be calibrators and $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$. For $\alpha\in(0,1)$, define the randomized rejection region by
$R_{\alpha}=\left\{(\mathbf{p},u)\in[0,1]^{K+1}:\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq u\right\}$
where we set $f_{k}(p_{k}/u)=0$ if $u=0$. Using $R_{\alpha}$, we can define the function $F:[0,1]^{K+1}\to[0,1]$ by
$F(\mathbf{p},u)$ $=\inf\{\alpha\in(0,1):(\mathbf{p},u)\in R_{\alpha}\}$
$=\inf\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq u\right\},$ (9)
with the convention $\inf\varnothing=1$ and $0\times\infty=\infty$ (this guarantees $F(\mathbf{p},u)=0$ when any component of $(\mathbf{p},u)$ is 00).
Theorem 3.12.
If $f_{1},\dots,f_{K}$ are calibrators and $(\lambda_{1},\dots,\lambda_{K})\in\Delta_{K}$, then $F$ in (3.3) is a homogenous randomized p-merging function. Moreover, $F$ is lower semicontinuous.
Proof.
It is clear that $F$ is increasing and Borel since $R_{\alpha}$ is a lower set. For $(\mathbf{P},U)=(P_{1},\dots,P_{K},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$ and $\alpha\in(0,1)$, using Theorem 3.10 and the fact that $(R_{\beta})_{\beta\in(0,1)}$ is nested, we have
$\mathbb{P}(F(\mathbf{P},U)\leq\alpha)$ $=\mathbb{P}\left(\inf\left\{\beta\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\beta}\right)\geq U\right\}\leq\alpha\right)$
$=\mathbb{P}\left(\bigcap_{\beta>\alpha}\left\{\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\beta}\right)\geq U\right\}\right)$
$=\inf_{\beta>\alpha}\mathbb{P}\left(\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{P_{k}}{\beta}\right)\geq U\right)\leq\inf_{\beta>\alpha}\beta=\alpha.$ (10)
Therefore, $F$ is a randomized p-merging function. Homogeneity of $F$ follows from (3.3).
Since $F$ is homogeneous and increasing, it is continuous in $\mathbf{p}$. Moreover, for fixed $\mathbf{p}\in[0,1]^{K}$, since
$\bigcap_{v<u}\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq v\right\}=\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq u\right\},$
we have
$\lim_{v\uparrow u}F(\mathbf{p},v)$ $=\lim_{v\uparrow u}\inf\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq v\right\}$
$=\inf\bigcap_{v<u}\left\{\alpha\in(0,1):\sum_{k=1}^{K}\lambda_{k}f_{k}\left(\frac{p_{k}}{\alpha}\right)\geq v\right\}=F(\mathbf{p},u).$
Therefore, $u\mapsto F(\mathbf{p},u)$ is lower semi-continuous. ∎
In case of symmetric p-merging functions (i.e., $F(\mathbf{p},u)=F(\mathbf{q},u)$ for any permutation $\mathbf{q}$ of $\mathbf{p}$), we have the following corollary, which directly follows from Theorem 3.12.
Corollary 3.13.
For any calibrator $f$, the function $F$ defined by
$F(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}f\left(\frac{p_{k}}{\alpha}\right)\geq u\right\},$ (11)
is a homogenous and symmetric randomized p-merging function.
A simple observation is that replacing $f(p)$ with $f(p)\wedge K$ does not change the function $F$. This observation allows us to only consider calibrators that are bounded above by $K$, which can improve some existing p-merging functions. This is similar to what was done in Vovk et al., (2022) in the context of deterministic p-merging functions.
Remark 3.14 (Internal randomization).
Note that $F$ in (3.3) is increasing in each of its arguments. If one has prior information that one of the p-values, say $P_{1}$, is independent of the rest (but the rest can be arbitrarily dependent), then one can use $P_{1}$ for randomization and apply $F$ (with one less input dimension for $\mathbf{p}$) to $(\mathbf{p},u)=(P_{2},\dots,P_{K},U)$ with $U=P_{1}$ to obtain a valid p-value that does not depend on external randomization. Increasing monotonicity of $u\mapsto F(\mathbf{p},u)$ guarantees two things. First, a p-variable may be stochastically larger than a standard uniform one, so increasing monotonicity is needed for validity. Second, if $P_{1}$ is indeed very small, i.e., it carries signal against the null, then the combined p-value will benefit from this signal. This form of internal randomization has been discussed in Wang, (2024, Section B.1). An alternative method of internal randomization through data (instead of p-values) is discussed by Ramdas and Manole, (2023, Section 10.6) but different from our setting.
It is feasible to combine the results presented in Sections 3.1 and 3.3 through the formulation of novel p-merging functions that exploit both the properties of exchangeability and randomization. These results are presented in Appendix B and are based on the exchangeable and uniformly randomized Markov inequality presented in Theorem 2.6.
3.4 Instantiating the above ideas
The ideas above were admittedly somewhat abstract, but provide us with the general tools to improve specific combination rules. The following sections do this for several rules, one by one. To elaborate, one of the most commonly employed p-merging functions is the Bonferroni method:
$F_{\mathrm{Bonf}}(\mathbf{p})=Kp_{(1)},$
where $p_{(1)}$ is the minimum of observed p-values. Rüger, (1978) extended the aforementioned rule in a more general sense. In particular, it is possible to prove that
$F_{\mathrm{R}}(\mathbf{p}):=\frac{K}{k}p_{(k)},\quad k\in\{1,\dots,K\},$ (12)
is a valid p-value, where $p_{(k)}$ represents the $k$-th smallest p-value among $(p_{1},\dots,p_{K})$. In other words, the $\lambda$-quantile $p_{(\lceil\lambda K\rceil)}$ is a valid p-value if multiplied by the factor $1/\lambda$. The next section improves on this combination rule.
Vovk and Wang, (2020) introduced the class of p-merging functions based on the generalized mean, also called M-family. This general class takes the form
$a_{r,K}\left(\frac{p_{1}^{r}+\dots+p_{K}^{r}}{K}\right)^{1/r},$ (13)
where $r\in\mathbb{R}\setminus\{0\}$ and $a_{r,K}$ is the smallest constant making (13) a p-merging function. This class encompasses numerous well-known cases, each distinguished by different values of the parameter $r$. In particular, if $r=1$ then (13) reduces to the simple average introduced in Rüschendorf, (1982) and the value $a_{1,K}=2$. Another important case is the harmonic mean obtained with $r=-1$. In the following sections, we demonstrate that if p-values exhibit exchangeability or if randomization is allowed, then it becomes feasible to enhance most of these combination rules.
3.5 Combining asymptotic p-values
Before proceeding with the remainder of the paper and introducing new merging rules based on the results presented in the preceding sections, we want to examine the scenario wherein the p-values are asymptotically valid. Many of the results obtained in the literature rely on uniform or super-uniform p-values (see Section 2); however, in statistical applications, p-values are often asymptotic, and they are not necessarily valid p-values in finite sample. See, for example, Severini, (2000, Chapter 4) for an introduction to p-values obtained using likelihood-based methods.
All methods in our paper work also for asymptotic p-values, i.e., those that converge in distribution to valid p-values. Suppose that $\mathbf{(}\mathbf{P}_{n})_{n\in\mathbb{N}}$ is a sequence of nonnegative random vectors that converges to a vector $\mathbf{P}$ of p-values in distribution. With an upper semicontinuous calibrator $f$ (recall that all admissible calibrators are upper semicontinuous), for each $\alpha\in(0,1)$ and $u\in(0,1)$, the rejection sets $R_{\alpha}$ given by
$R_{\alpha}=\left\{(p_{1},\dots,p_{K})\in[0,\infty)^{K}:\exists k\leq K\textrm{such that }\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\right\}$
or
$R_{\alpha}=\left\{(p_{1},\dots,p_{K})\in[0,\infty)^{K}:\frac{1}{K}\sum_{k=1}^{K}f\left(\frac{p_{k}}{\alpha}\right)\geq u\right\}$
are closed. As a consequence, by the Portmanteau Theorem,
$\limsup_{n\to\infty}\mathbb{P}(\mathbf{P}_{n}\in R_{\alpha})\leq\mathbb{P}(\mathbf{P}\in R_{\alpha}).$
Therefore, any methods in our paper that produce a valid p-value for the vector $\mathbf{P}$ of p-values (exchangeable or arbitrarily dependent) produce an asymptotic p-value for any $(\mathbf{P}_{n})_{n\in\mathbb{N}}$ that converges to $\mathbf{P}$ in distribution.
4 Improving Rüger’s combination rule
Vovk et al., (2022) showed that the p-merging function defined in (12), with a trivial modification (i.e., return 00 if $p_{(1)}=0$; see Theorem 7.3 of Vovk et al., (2022)) is admissible for $k\neq K$, and it is admissible among symmetric p-merging functions when $k=K$. In particular, the corresponding calibrator that induces (12) is
$f(p)=\frac{K}{k}\mathds{1}\{p\in(0,k/K]\}+\infty\mathds{1}\{p=0\},$
and this implies that we can exploit directly the duality between rejection regions and p-merging functions.
4.1 An exchangeable Rüger combination rule
If the exchangeability condition is satisfied, then we can obtain something better than the combination rule in (12). First, it is clear that $f(p)=\frac{K}{k}\mathds{1}\{p\in(0,k/K]\}+\infty\mathds{1}\{p=0\}$ is an admissible calibrator. We now define the function
$F_{\mathrm{ER}}(\mathbf{p}):=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\frac{K}{k}\mathds{1}\left\{\frac{p_{i}}{\alpha}\leq\frac{k}{K}\right\}\geq 1\right\}.$ (14)
Below, for fixed $k\in\{1,\dots,K\}$, we let $p^{\ell}_{(\lambda_{\ell})}$ denote the $\lceil\ell\frac{k}{K}\rceil$-th ordered value obtained using the first $\ell$ values of $\mathbf{p}$. Essentially, $p^{\ell}_{(\lambda_{\ell})}$ is the upper quantile of order $k/K$ obtained using the first $\ell$ p-values.
Theorem 4.1.
For any fixed $k\in\{1,\dots,K\}$ the function $F_{\mathrm{ER}}$ defined in (14) satisfies
$F_{\mathrm{ER}}(\mathbf{p})=\left(\frac{K}{k}\bigwedge_{\ell=1}^{K}p^{\ell}_{(\lambda_{\ell})}\right)\mathds{1}\{p_{(1)}>0\}\mbox{\leavevmode\nobreak\ for $\mathbf{p}\in[0,1]^{K}$, where $\lambda_{\ell}:=\lceil\ell\frac{k}{K}\rceil$,}$
and it is an exchangeable p-merging function that strictly dominates $F_{\mathrm{R}}$ in (12).
Proof.
According to Theorem 3.4, it follows that the function $F_{\mathrm{ER}}$ is a valid exchangeable p-merging function. Fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$. Note that $F_{\mathrm{ER}}(\mathbf{p})\leq\alpha$ if and only if
$\exists\ell\leq K:\frac{1}{\ell}\sum_{i=1}^{\ell}\frac{K}{k}\mathds{1}\left\{\frac{p_{i}}{\alpha}\leq\frac{k}{K}\right\}\geq 1\implies\exists\ell\leq K:\sum_{i=1}^{\ell}\mathds{1}\left\{p_{i}\leq\alpha\frac{k}{K}\right\}\geq\left\lceil\ell\frac{k}{K}\right\rceil.$
Where rounding up is due to the fact that the summation takes values in positive integers. This holds true if and only if
$\exists\ell\leq K:p^{\ell}_{(\lambda_{\ell})}\leq\alpha\frac{k}{K},$
where $p^{\ell}_{(\lambda_{\ell})}$ is the $\lceil\ell\frac{k}{K}\rceil$-th ordered value of $(p_{1},\dots,p_{\ell})$. Rearranging the terms, we obtain that it is verified when
$\frac{K}{k}\bigwedge_{\ell=1}^{K}p_{(\lambda_{\ell})}^{\ell}\leq\alpha,$
which complete the first part of the proof. For the second statement, it is possible to note that the element $p_{(\lambda_{\ell})}^{\ell}$ in the sequence coincides with $p_{(k)}$ when $\ell=K$. ∎
It may be useful to note that the Bonferroni rule is not improved using this method. Indeed, fixing $k=1$, we find that $p_{(\lambda_{\ell})}^{\ell}$ reduces to the minimum of the first $\ell$ p-values subsequently taking the minimum of the obtained sequences coincides with the overall minimum. In addition, Rüger can be sharp, for some exchangeable $\mathbf{P}$, i.e., satisfying $F_{\rm R}(\mathbf{P})\in\mathcal{U}$, and so is our proposal.
4.2 A randomized Rüger combination rule
In this part, we prove that if randomization is allowed, it becomes feasible to enhance the combination introduced by Rüger, (1978), even if the sequence of p-values presents an arbitrary dependence and the obtained result has nice properties in terms of interpretability. As before, we define the merging function
$F_{\mathrm{UR}}(\mathbf{p},u):=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\frac{K}{k}\mathds{1}\left\{\frac{p_{k}}{\alpha}\leq\frac{k}{K}\right\}\geq u\right\}.$ (15)
Theorem 4.2.
For any fixed $k\in\{1,\dots,K\}$, the function $F_{\mathrm{UR}}$ from (15) satisfies
[…]
and it is a valid randomized p-merging function that strictly dominates $F_{\mathrm{R}}$ in (12).
Proof.
According to Corollary 3.13, it follows that the function $F_{\mathrm{UR}}$ is a valid randomized p-merging function. Fix any $\alpha\in(0,1)$ and $(\mathbf{p},u)\in(0,1]^{K+1}$, then it is possible to note that $F(\mathbf{p},u)\leq\alpha$ if and only if
$\frac{1}{K}\sum_{i=1}^{K}\frac{K}{k}\mathds{1}\left\{\frac{p_{i}}{\alpha}\leq\frac{k}{K}\right\}\geq u\implies\sum_{i=1}^{K}\mathds{1}\left\{p_{i}\leq\alpha\frac{k}{K}\right\}\geq\lceil uk\rceil.$
Rearranging the terms this holds true only if
$p_{(\lceil uk\rceil)}\leq\alpha\frac{k}{K}\implies\frac{K}{k}p_{(\lceil uk\rceil)}\leq\alpha,$
which concludes the claim. Since $u\leq 1$ almost surely, we have $p_{(\lceil uk\rceil)}\leq p_{(k)}$. ∎
The above theorem implies in particular that the Rüger combination rule is inadmissible if external randomization is allowed, despite it being admissible if randomization is not allowed (Vovk et al.,, 2022). The fact that $p_{(\lceil uK\rceil)}$ is a valid p-value is particularly interesting. It has a simple interpretation: sort the p-values and pick the one at a uniformly random index. In addition, for the latter combination rule (8) holds for all $\beta\in(0,1]$.
It is worth noting that when $k=1$ the Rüger combination rule reduces to the Bonferroni method; however, the introduction of a randomization does not yield any practical benefit since $\lceil u\rceil=1$.
5 Improving Hommel’s combination rule
The method proposed in Section 4 requires us to choose the value of $k$ in advance; a solution that solves the problem is proposed in Hommel, (1983). The combination rule, in this case, is given by
$F^{\prime}_{\mathrm{Hom}}(\mathbf{p}):=h_{K}\bigwedge_{k=1}^{K}F_{\mathrm{R}}(\mathbf{p};k)=\left(\sum_{k=1}^{K}\frac{1}{k}\right)\bigwedge_{k=1}^{K}\frac{K}{k}p_{(k)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{with }h_{K}=\sum_{k=1}^{K}\frac{1}{k}.$ (16)
Essentially, this function allows selecting the minimum derived from the combinations based on ordered statistics with the additional cost of a factor equal to $h_{K}\approx\log K$.
It is possible to prove that the Hommel combination rule is not admissible and is dominated by the grid harmonic merging function introduced in Vovk et al., (2022). For completeness, we state here a useful lemma.
Lemma 5.1.
Let $f$ be a function defined by
$f(p)=\frac{K\mathds{1}\{h_{K}p\leq 1\}}{\lceil Kh_{K}p\rceil}.$ (17)
Then $f$ is an admissible calibrator. Moreover, the p-merging function induced by $f$ is
$F_{\mathrm{Hom}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\sum_{k=1}^{K}\frac{\mathds{1}\{h_{K}p_{k}/\alpha\leq 1\}}{\lceil Kh_{K}p_{k}/\alpha\rceil}\geq 1\right\},$ (18)
is valid and it dominates the Hommel combination rule.
The proof is given in Appendix A. The calibrator in (17) coincides, with a slight adjustment, with the BY calibrator introduced in Xu et al., (2024). An interesting fact is that the function $F_{\mathrm{Hom}}$ is always admissible in the class of symmetric p-merging functions; while it is admissible in the class of p-merging function (not necessarily symmetric) if $K$ is not a prime number (Vovk et al.,, 2022, Theorem 7.1). The Hommel function allows for the selection of the minimum among the $K$ possible different quantiles of $\mathbf{p}$.
In reality, one can choose to select only certain quantiles among $K$ (e.g., one can select the minimum between $K$ times the minimum, 2 times the median and the maximum), hoping to pay a price less than $\log K$. We treat this problem in Appendix C, where we introduce a generalization of the Hommel combination rule.
5.1 An exchangeable Hommel’s combination rule
Starting from the results in the previous sections, we can introduce a merging function, which improves Hommel’s combination if the input p-values are exchangeable. In particular, we define the function
$F_{\mathrm{EHom}}(\mathbf{p}):=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\frac{K\mathds{1}\{h_{K}p_{i}/\alpha\leq 1\}}{\lceil Kh_{K}p_{i}/\alpha\rceil}\geq 1\right\},$ (19)
which can be considered an improvement of the function $F_{\mathrm{Hom}}$ defined in (18).
Theorem 5.2.
The function $F_{\mathrm{EHom}}$ defined in (19) is a valid exchangeable p-merging function and it strictly dominates the function $F_{\mathrm{Hom}}$ defined in (18).
Proof.
According to Theorem 3.4 and Lemma 5.1, it follows that $F_{\mathrm{EHom}}$ is a valid exchangeable p-merging function. It is simple to see that $F_{\mathrm{EHom}}\leq F_{\mathrm{Hom}}$. ∎
The computation of a closed form for (19) is not straightforward, a possible solution to calculate the value of $F_{\mathrm{EHom}}$ is by using Algorithm 1 defined in Appendix E.
5.2 A randomized Hommel’s combination rule
The randomized version of the function $F_{\mathrm{Hom}}$ in (18) has been proposed in Xu and Ramdas, (2023, Appendix E) and takes the following form:
$F_{\mathrm{UHom}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\sum_{k=1}^{K}\frac{\mathds{1}\{h_{K}p_{k}/\alpha\leq 1\}}{\lceil Kh_{K}p_{k}/\alpha\rceil}\geq u\right\}.$ (20)
For completeness, we report here the following theorem.
Theorem 5.3.
The function $F_{\mathrm{UHom}}$ defined in (20) is a valid p-merging function and it strictly dominates the function $F_{\mathrm{Hom}}$ defined in (18).
Proof.
According to Corollary 3.13 and Lemma 5.1, it follows that $F_{\mathrm{UHom}}$ is a valid randomized p-merging function. It is simple to see that $F_{\mathrm{UHom}}\leq F_{\mathrm{Hom}}$. ∎
The value of function (20) can be computed using Algorithm 1 in Vovk et al., (2022), substituting the value of $1$ by $u$.
6 Improving the “twice the average” combination rule
We now study the case of $r=1$ in (13), which corresponds to the arithmetic mean. The general case, which allows $r\in\mathbb{R}\setminus\{0\}$, is considered in Appendix D. In the following, let $A(\mathbf{p})$ denote the arithmetic average of any vector $\mathbf{p}$, let $\mathbf{p}_{m}:=(p_{1},\dots,p_{m})$ denote the vector containing the first $m$ p-values, and let $\mathbf{p}_{(m)}$ denote the vector containing the smallest $m$ elements of $\mathbf{p}$: $\mathbf{p}_{(m)}=(p_{(1)},\dots,p_{(m)})$ such that $p_{(1)}\leq\dots\leq p_{(m)}$. In addition, we denote by $\mathbf{p}_{(m)}^{\ell}=(p_{(1)}^{\ell},\dots,p_{(m)}^{\ell})$, $m\in\{1,\dots,\ell\}$, the vector containing the smallest $m$ elements of $(p_{1},\dots,p_{\ell})$. First, let us introduce a lemma that will be instrumental in subsequent discussions.
Lemma 6.1.
Let $f(p)=(2-2p)_{+}\mathds{1}\{p\in(0,1]\}+\infty\mathds{1}\{p=0\}$. Then, $f$ is an admissible calibrator.
Proof.
It is easy to see that $f$ is decreasing and $\int_{0}^{1}(2-2p)\mathrm{d}p=1.$ In addition, it is upper semi-continuous and $f(0)=\infty$. ∎
In particular, we have that the calibrator defined in Lemma 6.1 is larger or equal than $f^{\prime}(p):=(2-2p)$, which is the function inducing the average combination rule, that we will denote
$F^{\prime}_{\mathrm{A}}(\mathbf{p}):=2A(\mathrm{p}),$ (21)
that is a valid p-merging function. Note that $f^{\prime}$ can take negative values, so it is technically not a calibrator in the sense of our definitions, since it can take in arguments larger than $1$ (like $p/\alpha$).
6.1 Exchangeable average combination rule
We now define a homogeneous p-merging function in this way
$F_{\mathrm{EA}}(\mathbf{p}):=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(2-2\frac{p_{i}}{\alpha}\right)_{+}\geq 1\right\}.$ (22)
Further, if we define the function
$F_{\mathrm{EA}}^{\prime}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(2-2\frac{p_{i}}{\alpha}\right)\geq 1\right\}$ (23)
then it holds that $F_{\mathrm{EA}}(\mathbf{p})\leq F_{\mathrm{EA}}^{\prime}(\mathbf{p})$. In particular, it is possible to note that $F_{\mathrm{EA}}^{\prime}(\mathbf{p})\leq\alpha$ if there exists at least one $\ell\in\{1,\dots,K\}$ such that $2\ell^{-1}\sum_{i=1}^{\ell}p_{i}\leq\alpha$.
Theorem 6.2.
The function $F^{\prime}_{\mathrm{EA}}$ defined in (23) equals
$2\left\{\bigwedge_{m=1}^{K}A(\mathbf{p}_{m})\right\}\,,$
and is a valid exchangeable p-merging function that strictly dominates $F_{\mathrm{A}}^{\prime}(\mathbf{p})$ from (21). However, it is itself strictly dominated by the function $F_{\mathrm{EA}}$ defined in (22), which equals
[…]
and is also a valid exchangeable p-merging function.
It is worth remarking that despite being strictly dominated, $F^{\prime}_{\mathrm{EA}}$ is very interpretable: it is just the minimum (over $m$) of “twice the average” of the first $m$ p-values.
Proof.
According to Theorem 3.4 and Lemma 6.1, it follows that the function $F_{\mathrm{EA}}$ is a valid exchangeable p-merging function. Fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$. We note that $F_{\mathrm{EA}}(\mathbf{p})\leq\alpha$ if and only if, for some $\ell\in\{1,\dots,K\}$, we have
$\frac{1}{\ell}\sum_{i=1}^{\ell}\left(2-\frac{2p_{i}}{\alpha}\right)_{+}\geq 1.$ (24)
This implies that there exists $\ell\leq K$ such that
$\frac{1}{\ell}\sum_{j=1}^{m}\left(2-\frac{2p^{\ell}_{(j)}}{\alpha}\right)\geq 1\mathrm{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\},$
where we recall that $p^{\ell}_{(j)}$ is the $j$-th ordered value of the vector $(p_{1},\dots,p_{\ell})$. This is due to the fact that the contribution of $p_{i}$ in the left-hand side of (24) vanishes for large values of $p_{i}$. Rearranging the terms, it is possible to obtain that exists $\ell\leq K$ such that
$\frac{2A(\mathbf{p}^{\ell}_{(m)})}{2-\ell/m}\leq\alpha\mathrm{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\}.$
Taking an infimum over $m$ yields
$\left(\bigwedge_{m=1}^{\ell}\frac{2A(\mathbf{p}^{\ell}_{(m)})}{(2-\ell/m)_{+}}\right)\leq\alpha\mathrm{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ some}\leavevmode\nobreak\ \ell\in\{1,\dots,K\}.$
Actually, the index $m$ can start from $\lceil\ell/2\rceil$ since the first $\lceil\ell/2\rceil-1$ terms in the denominator are smaller than zero. Taking an infimum also over $\ell$ gives the desired result. ∎
6.2 Randomized average combination rule
We now derive an improvement for the “twice the average” rule using a simple randomization trick. In this case, we do not require exchangeability but we allow for an arbitrary dependence among the p-values. We define the randomized p-merging function $F_{\mathrm{UA}}$ by
$F_{\mathrm{UA}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(2-\frac{2p_{k}}{\alpha}\right)_{+}\geq u\right\}.$ (25)
Clearly, $F_{\mathrm{UA}}(\mathbf{p},u)\leq F^{\prime}_{\mathrm{UA}}(\mathbf{p},u)$, where $F^{\prime}_{\mathrm{UA}}$ is defined by
$F^{\prime}_{\mathrm{UA}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(2-\frac{2p_{k}}{\alpha}\right)\geq u\right\}=\frac{2A(\mathbf{p})}{2-u}.$ (26)
In particular, if randomization is not allowed then $u$ is replaced by $1$ and $F^{\prime}_{\mathrm{UA}}(\mathbf{p},1)$ coincides with the standard 2 times the average of p-value described in (21), and it is dominated by
$F_{A}(\mathbf{p})=\left(\bigwedge_{m=1}^{K}\frac{2A(\mathbf{p}_{(m)})}{(2-K/m)_{+}}\right)\mathds{1}\{p_{(1)}>0\},$
that is the non-randomized version of $F_{\mathrm{UA}}$. Now, we obtain a simple formula for $F_{\mathrm{UA}}$.
Theorem 6.3.
The function $F^{\prime}_{\mathrm{UA}}$ defined in (26) equals
$\frac{2A(\mathbf{p})}{2-u},$
is a valid randomized p-merging function, and it strictly dominates the function $F_{\mathrm{A}}^{\prime}(\mathbf{p})$ from (21). However, it is strictly dominated by $F_{\mathrm{UA}}$ in (25), which equals
$\left(\bigwedge_{m=1}^{K}\frac{2A(\mathbf{p}_{(m)})}{(2-Ku/m)_{+}}\right)\mathds{1}\{p_{(1)}>0\}$
and is also a valid randomized p-merging function.
Proof.
According to Corollary 3.13 and Lemma 6.1, it follows that the function $F_{\mathrm{UA}}$ is a valid randomized p-merging function. Fix any $\alpha\in(0,1)$ and $(\mathbf{p},u)\in(0,1]^{K+1}$. Note that $F_{\mathrm{UA}}(\mathbf{p},u)\leq\alpha$ if and only if
$\frac{1}{K}\sum_{k=1}^{m}\left(2-\frac{2p_{(k)}}{\alpha}\right)\geq u\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some $m\in\{1,\dots,K\}$}.$
Rearranging terms, it is
$\frac{\sum_{k=1}^{m}{2p_{(k)}}}{2m-Ku}\leq\alpha\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some $m\in\{1,\dots,K\}$}.$
Taking an infimum over $m$ yields the desired formula. ∎
A method that can be directly compared with Theorem 6.3 is to use $F_{\mathrm{UA}}^{*}(\mathbf{p},u):=A(\mathbf{p})/(2-2u)$ proposed by Wang, (2024, Section B.2). This function $F_{\mathrm{UA}}^{*}$ is also a valid randomized p-merging function. One can see that $F_{\mathrm{UA}}^{*}$ does not dominate and is not dominated by any of $F_{\mathrm{A}}$, $F_{\mathrm{UA}}$ and $F_{\mathrm{UA}}^{\prime}$. Moreover, there is a simple relationship: $F_{\mathrm{UA}}^{\prime}(\mathbf{p},u)\leq F_{\mathrm{UA}}^{*}(\mathbf{p},u)$ if and only if $u\geq 2/3$ for every $\mathbf{p}$ that is not the zero vector.
7 Improving the harmonic mean combination rule
The harmonic mean p-value was studied by Wilson, (2019). In our context, it corresponds to the merging function in (13) when $r=-1$. We first state a lemma on a calibrator that we will use later.
Lemma 7.1.
Define the function
$f(p)=\min\left\{\frac{1}{T_{K}p}-\frac{1}{T_{K}},K\right\}\mathds{1}\{p\in[0,1]\},$
with $T_{K}\geq 1$. Then $f$ is a calibrator if $T_{K}$ is such that $KT_{K}+1-e^{T_{K}}\leq 0$, and in particular, $f$ is a calibrator if $T_{K}=\log K+\log\log K+1$.
The proof is given in Appendix A. In the following, we fix
$T_{K}=\log K+\log\log K+1$
and we denote by $H(\mathbf{p})=K(\sum_{k=1}^{K}1/p_{k})^{-1}$ the harmonic mean of the vector $\mathbf{p}$ where $K$ is the number of elements contained in $\mathbf{p}$. We begin with a result that is new even under arbitrary dependence.
Proposition 7.2.
$F^{\prime}_{\mathrm{H}}(\mathbf{p}):=(T_{K}+1)H(\mathbf{p})$ is a valid p-merging function under arbitrary dependence.
The proof is given in Appendix A. The above result differs from the formulation given in Vovk and Wang, (2020, Proposition 9), which states that $e\log KH(\mathbf{p})$ is a valid p-merging function, thus sharpening their result for $K\geq 4$. We will now improve this result for exchangeable p-values.
7.1 Exchangeable harmonic mean combination rule
We define an exchangeable homogeneous p-merging function as follows
$F_{\mathrm{EH}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(\frac{\alpha}{T_{K}p_{i}}-\frac{1}{T_{K}}\right)_{+}\geq 1\right\}.$ (27)
In addition, if we define the function
$F^{\prime}_{\mathrm{EH}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(\frac{\alpha}{T_{K}p_{i}}-\frac{1}{T_{K}}\right)\geq 1\right\},$ (28)
then it holds that $F_{\mathrm{EH}}(\mathbf{p})\leq F^{\prime}_{\mathrm{EH}}(\mathbf{p})$. In particular, it is possible to note that $F^{\prime}_{\mathrm{EH}}(\mathbf{p})\leq\alpha$ if $(T_{K}+1)H(\mathbf{p}_{\ell})\leq\alpha$ is true for some $\ell\in\{1,\dots,K\}$.
Theorem 7.3.
The function $F^{\prime}_{\mathrm{EH}}$ defined in (28) equals
$\bigwedge_{m=1}^{K}(T_{K}+1)H(\mathbf{p}_{m})$
and it is a valid exchangeable p-merging function. It strictly dominates $F^{\prime}_{\mathrm{H}}(\mathbf{p})$ in Proposition 7.2. However, it is itself strictly dominated by $F_{\mathrm{EH}}$ in (27), which equals
$\bigwedge_{\ell=1}^{K}\left(\bigwedge_{m=1}^{\ell}\left(\frac{\ell\,T_{K}}{m}+1\right)H(\mathbf{p}_{(m)}^{\ell})\right)$
and is also a valid exchangeable p-merging function.
Proof.
According to Theorem 3.4 and Lemma 7.1, it follows that $F_{\mathrm{EH}}$ is a valid exchangeable p-merging function. Fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$. We note that $F_{\mathrm{EH}}(\mathbf{p})\leq\alpha$ if and only if, for some $\ell\in\{1,\dots,K\}$, we have
$\frac{1}{\ell}\sum_{i=1}^{\ell}\left(\frac{\alpha}{T_{K}p_{i}}-\frac{1}{T_{K}}\right)_{+}\geq 1.$
This implies that exists $\ell\leq K$ such that
$\frac{1}{\ell}\sum_{j=1}^{m}\left(\frac{\alpha}{T_{K}p^{\ell}_{(j)}}-\frac{1}{T_{K}}\right)\geq 1\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\},$
where we recall that $p^{\ell}_{(j)}$ is the $j$-th ordered value of the vector $(p_{1},\dots,p_{\ell})$. Rearranging the terms it is possible to obtain that exist $\ell\leq K$ such that
$\left(\frac{\ell\,T_{K}}{m}+1\right)H(\mathbf{p}^{\ell}_{(m)})\leq\alpha\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\},$
Taking an infimum over $m$, we get
$\bigwedge_{m=1}^{\ell}\left(\frac{\ell\,T_{K}}{m}+1\right)H(\mathbf{p}^{\ell}_{(m)})\leq\alpha\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }\ell\in\{1,\dots,K\}.$
Taking an infimum over $\ell$ yields the desired result. ∎
7.2 Randomized harmonic mean combination rule
Similarly to Section 6.2, we derive an improvement for the harmonic mean using a randomization trick in the case of arbitrarily dependent p-values. We define the randomized p-merging function $F_{\mathrm{UH}}$ as follows
$F_{\mathrm{UH}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\alpha}{T_{K}p_{k}}-\frac{1}{T_{K}}\right)_{+}\geq u\right\}.$ (29)
It is possible to prove that $F_{\mathrm{UH}}(\mathbf{p},u)\leq F^{\prime}_{\mathrm{UH}}(\mathbf{p},u)$; where $F^{\prime}_{\mathrm{UH}}$ is
$F^{\prime}_{\mathrm{UH}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\alpha}{T_{K}p_{k}}-\frac{1}{T_{K}}\right)\geq u\right\}=(T_{K}u+1)H(\mathbf{p}).$ (30)
Theorem 7.4.
The function $F^{\prime}_{\mathrm{UH}}$ defined in (30) equals
$(T_{K}u+1)H(\mathbf{p}),$
and it is a valid randomized p-merging function. It strictly dominates the function $F^{\prime}_{\mathrm{H}}(\mathbf{p})$ in Proposition 7.2. However, it is itself strictly dominated by the function $F_{\mathrm{UH}}$ defined in (29), which equals
$\bigwedge_{m=1}^{K}\left(\frac{uKT_{K}}{m}+1\right)H(\mathbf{p}_{(m)})$
and is also a valid randomized p-merging function.
Proof.
According to Corollary 3.13 and Lemma 7.1, it follows that $F_{\mathrm{UH}}$ is a valid randomized p-merging function. Fix any $\alpha\in(0,1)$ and $(\mathbf{p},u)\in(0,1]^{K+1}$. Then $F_{\mathrm{UH}}(\mathbf{p},u)\leq\alpha$ if and only if
$\frac{1}{K}\sum_{k=1}^{m}\left(\frac{\alpha}{T_{K}p_{(k)}}-\frac{1}{T_{K}}\right)\geq u\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,K\}.$
Rearranging the terms, it is
$\left({uKT_{K}}+m\right)\left(\sum_{k=1}^{m}\frac{1}{p_{(k)}}\right)^{-1}\leq\alpha\quad\mathrm{for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,K\}.$
Taking a minimum over $m$ yields the desired formula. ∎
A non-randomized improvement of $F^{\prime}_{\mathrm{H}}$ can be achieved fixing $u=1$ in (29). This coincides with the function $F_{\mathrm{H}}(\mathbf{p})=\bigwedge_{m=1}^{K}((KT_{K})/m+1)H(\mathbf{p}_{(m)})$.
8 Improving the geometric mean combination rule
We now derive some new combination based on the geometric mean. In particular, the geometric mean can be studied as a special case of (13) when $r\to 0$. Let $G(\mathbf{p})=(\prod_{k=1}^{K}p_{k})^{1/k}$ denote the geometric mean of the vector $\mathbf{p}$. The calibrator, in this case, is given by $f(p)=(-\log p)_{+}$, which is an admissible calibrator. Actually, a slightly improved calibrator is $f(p)=(-(\log p)/T)_{+}\wedge K$ for some $T<1$ satisfying $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$. This condition is verified when $1-e^{-KT}\leq T$, which makes $T$ very close to 1 for $K$ moderately large (see Section 3.2 in Vovk and Wang,, 2020). In the sequel, we denote by
$F^{\prime}_{\mathrm{G}}(\mathbf{p}):=eG(\mathbf{p}),$ (31)
which is a valid p-merging function.
8.1 Exchangeable geometric mean combination rule
Following the same approach as in the preceding sections, we define the function $F_{\mathrm{EG}}$ as follows
$F_{\mathrm{EG}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}(-\log p_{i}+\log\alpha)_{+}\geq 1\right\}.$ (32)
The counterpart without truncation to positive values, $F_{\mathrm{EG}}$, is defined as
$F^{\prime}_{\mathrm{EG}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}(-\log p_{i}+\log\alpha)\geq 1\right\},$ (33)
and it holds that $F_{\mathrm{EG}}(\mathbf{p})\leq F^{\prime}_{\mathrm{EG}}(\mathbf{p})$.
Theorem 8.1.
The function $F^{\prime}_{\mathrm{EG}}$ defined in (33) equals
$e\left\{\bigwedge_{m=1}^{K}G(\mathbf{p}_{m})\right\}\,,$
is a valid exchangeable p-merging function, and it strictly dominates the function $F_{\mathrm{G}}(\mathbf{p})$ in (31). However, it is strictly dominated by the function $F_{\mathrm{EG}}$ defined in (32), which equals
$\bigwedge_{\ell=1}^{K}\left(\bigwedge_{m=1}^{\ell}e^{\ell/m}G(\mathbf{p}^{\ell}_{(m)})\right)$
and is also a valid exchangeable p-merging function.
Proof.
According to Theorem 3.4, it follows that the function $F_{\mathrm{EG}}$ is a valid exchangeable p-merging function. Fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$. Then $F_{\mathrm{EG}}(\mathbf{p})\leq\alpha$, if and only if exists $\ell\leq K$ such that
$\frac{1}{\ell}\sum_{i=1}^{\ell}(-\log p_{i}+\log\alpha)_{+}\geq 1.$
This is verified when exists $\ell\leq K$ such that
$\frac{1}{\ell}\sum_{j=1}^{m}(-\log p^{\ell}_{(j)}+\log\alpha)\geq 1\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\},$
where we recall that $p^{\ell}_{(j)}$ is the $j$-th ordered value of the vector $(p_{1},\dots,p_{\ell}).$ Rearranging the terms it is possible to obtain that exists $\ell\leq K$ such that
$e^{\ell/m}G(\mathbf{p}^{\ell}_{(m)})\leq\alpha\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,\ell\}.$
Taking an infimum over $m$, we get
$\bigwedge_{m=1}^{\ell}\left(e^{\ell/m}G(\mathbf{p}^{\ell}_{(m)})\right)\leq\alpha\quad\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }\ell\in\{1,\dots,K\}.$
Taking an infimum over $\ell$ yields the desired result. ∎
8.2 Randomized geometric mean combination rule
As in the previous sections, we define the randomized p-merging function $F_{\mathrm{UG}}$ as follows:
$F_{\mathrm{UG}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}(-\log p_{k}+\log\alpha)_{+}\geq u\right\}.$ (34)
In this situation, we can define the function $F_{\mathrm{UG}}^{\prime}$ by
$F^{\prime}_{\mathrm{UG}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}(-\log p_{k}+\log\alpha)\geq u\right\}=e^{u}G(\mathbf{p}).$ (35)
It is simple to prove that $F_{\mathrm{UG}}(\mathbf{p},u)\leq F^{\prime}_{\mathrm{UG}}(\mathbf{p},u)$. The function $F^{\prime}_{\mathrm{UG}}(\mathbf{p},u)$ can be seen as a randomized improvement over $eG(\mathbf{p})$, a deterministic geometric p-merging function studied by Vovk and Wang, (2020).
Theorem 8.2.
The function $F^{\prime}_{\mathrm{UG}}$ defined in (35) equals
$e^{u}G(\mathbf{p}),$
is a valid randomized p-merging function, and it strictly dominates the function $F^{\prime}_{\mathrm{G}}(\mathbf{p})$ in (31). However, it is strictly dominated by $F_{\mathrm{UG}}$ in (34), which equals
$\bigwedge_{m=1}^{K}\left(e^{u\frac{K}{m}}G(\mathbf{p}_{(m)})\right)$
and is also a valid randomized p-merging function.
Proof.
According to Corollary 3.13 and Lemma 7.1, it follows that $F_{\mathrm{UG}}$ is a valid randomized p-merging function. Fix any $\alpha\in(0,1)$ and $(\mathbf{p},u)\in(0,1]^{K+1}$. We have that $F_{\mathrm{UG}}(\mathbf{p},u)\leq\alpha$ if and only if
$\frac{1}{K}\sum_{k=1}^{m}(-\log p_{(k)}+\log\alpha)\geq u\quad\mathrm{for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,K\}.$
Rearranging the terms, it is
$e^{u\frac{K}{m}}G(\mathbf{p}_{(m)})\leq\alpha\quad\mathrm{for\leavevmode\nobreak\ some\leavevmode\nobreak\ }m\in\{1,\dots,K\}.$
Taking a minimum over $m$ yields the desired formula. ∎
A non-randomized improvement of the combination in (31) can be obtained fixing $u=1$ in (34). This gives the combination rule $F_{\mathrm{G}}(\mathbf{p})=\bigwedge_{m=1}^{K}e^{K/m}G(\mathbf{p}_{(m)})$.
9 Simulation study
In the previous few sections, new p-merging functions have been introduced. These new rules are obtained using a randomization trick (Section 3.3) or they rely on exchangeability of p-values (Section 3.1). Specifically, the introduced rules have been shown to dominate their original counterparts by utilizing randomness or exchangeability (or both). In this section, our aim is to investigate their performance using simulated data. Performance will be evaluated both in terms of power relative to their original counterparts and among various families of p-merging functions.
Figure 1: Combination of p-values using different rules. Every subplot illustrates power against $\mu$. The left endpoint of $\mu=0$ actually represents the empirical type-I error, which is controlled at the nominal level $\alpha=0.05$ for all methods proposed. The top row has $\rho=0.9$, while the bottom row has $\rho=0.1$ — as expected, the Bonferroni correction is more powerful near independence, but is less powerful under strong dependence. Further, the exchangeable and randomized improvements offer sizeable increases in power over the original variants in all settings.
We consider the example described in Vovk and Wang, (2020, Section 6) (a similar example is proposed in Chen et al., (2023)), where p-values are generated in the following way:
$X_{k}=\rho Z+\sqrt{1-\rho^{2}}Z_{k}-\mu,\quad p_{k}=\Phi(X_{k}),$ (36)
where $\Phi(\cdot)$ is the cumulative density function of the standard normal distribution, $Z,Z_{1},\dots,Z_{K}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$, and $\mu\geq 0$ and $\rho\in[0,1]$ are constants. It is simple to prove that $p_{1},\dots,p_{K}$ are exchangeable and their marginal distribution does not depend on $\rho$. In addition, if $\rho=0$ then $p_{1},\dots,p_{K}$ are independent while if $\rho=1$ then $p_{1}=\dots=p_{K}$. The value $p_{k}$, in this case, can be interpreted as the p-value resulting from a one-side z-test of the null hypothesis $\mu=0$ against the alternative $\mu>0$ from the statistic $X_{k}\sim\mathcal{N}(-\mu,1)$ with unknown $\mu$.
We let the parameter $\mu$ vary in the interval $[0,3]$ (if $\mu=0$ then $H_{0}$ is true) and fix the upper bound of the type I error to the nominal level $0.05$. We test the merging rules introduced in Section 4 and in Section 6, specifically the rules: “twice the median” and “twice the mean”. The parameter $\rho=\{0.1,0.9\}$, corresponding to weak and strong dependence among p-values. Each simulation is repeated for a total of $B=10,000$ replications, and we report the observed empirical average.
In Figure 1 we can notice that the nominal error level is controlled at the nominal level $0.05$ for all proposed methods. In the case of the Rüger combination, in both dependency scenarios, the power of the combinations obtained by exploiting exchangeability or employing external randomization is highly similar. In the case of the combination based on the arithmetic mean, it appears that the rules obtained using randomization exhibit greater power than those based on exchangeability (and, naturally, than the original rules). Additionally, in the case of the “twice the mean” rule, the advantages of using the rule obtained utilizing the calibrator $(2-2p)_{+}$ compared to the standard rule can be observed (omitted here; see Appendix F). In general, across all observed scenarios, the new rules demonstrate a quite significant improvement in terms of power.
In Figure 2, we compare different exchangeable p-merging functions introduced in the previous sections, with the addition of the Bonferroni method. P-values are generated as in (36), with $\rho=\{0.9,0.1\}$ and $\mu\in[0,3]$. The parameter $k$ for the randomized Rüger combination rule is set to $K/2$. The value of $\alpha$ is set to 0.05 and each simulation is repeated for a total of $B=10,000$ replications in which the observed empirical average is reported. In Figure 2, variations in terms of power are observed depending on whether the parameter $\rho$ is set to 0.9 or 0.1. Specifically, exchangeable p-merging functions that exhibit strong performance in the left plot tend to show reduced effectiveness in the right plot, and the opposite is also true. As expected, the Bonferroni method shows a higher power near independence where also the exchangeable Hommel combination rule defined in (19) seems to perform well.
Figure 2: Combination of p-values using different exchangeable p-merging functions. The performance of the different exchangeable p-merging functions is almost reversed in the two situations.
10 Summary
In this paper, we establish novel p-merging functions utilized in the scenario where the sequence of p-values is exchangeable. These new rules are demonstrated to dominate their original counterparts derived under the assumption of arbitrary dependence. Furthermore, we illustrate how a simple randomization trick (introduction of a uniform random variable or uniform permutation) can also be employed in the case of arbitrarily dependent p-values to yield more powerful rules than existing ones. These results are proposed in a fully general form, addressing the relationship between p-merging functions and e-values introduced by their respective calibrators. In particular, once the corresponding e-value is obtained, it becomes feasible to utilize Markov’s inequality or its randomized and exchangeable generalizations from Ramdas and Manole, (2023). Various subcases are explored individually, notably the quantile-based combination and some instances within the M-family.
Acknowledgments
The authors thank Ilmun Kim for a suggestion on asymptotic p-values.
Appendix A Omitted proofs
Proof of Lemma 5.1.
It is simple to see that $f$ is decreasing and it is upper semicontinuous (the function $f$ has discontinuity points in $i/(Kh_{K}),\leavevmode\nobreak\ i=1,\dots,K,$ but it is simple to prove that $\lim_{x\to x_{0}}f(x)\leq f(x_{0})$). In addition,
$\int_{0}^{1}f(p)\mathrm{d}p=\int_{0}^{1}\frac{K\mathds{1}\{h_{K}p\leq 1\}}{\lceil Kh_{K}p\rceil}\mathrm{d}p=\sum_{j=1}^{K}\frac{K}{j}\frac{1}{Kh_{K}}=\frac{1}{h_{K}}\sum_{j=1}^{K}\frac{1}{j}=1.$
Due to Theorem 2.2 we have that $F_{\mathrm{Hom}}$ is admissible.
To prove the last part, we see that for any $\mathbf{p}\in(0,1]^{K}$ and $\alpha\in(0,1)$ we have that $F_{\mathrm{Hom}}(\mathbf{p})\leq\alpha$, if and only if
$\bigwedge_{k=1}^{K}\frac{K}{k}p_{(k)}\leq\frac{\alpha}{h_{K}}.$
This implies that, for some $m\in\{1,\dots,K\}$, we have that
$\sum_{j=1}^{K}\mathds{1}\left\{\frac{K}{m}h_{K}p_{j}\leq\alpha\right\}\geq m.$
We now define this chain of inequalities,
$1\leq\sum_{j=1}^{K}\frac{1}{m}\mathds{1}\left\{\frac{K}{m}h_{K}p_{j}\leq\alpha\right\}\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{j=1}^{K}\frac{1}{\lceil Kh_{K}p_{j}/\alpha\rceil}\mathds{1}\left\{\frac{K}{m}h_{K}p_{j}\leq\alpha\right\}\stackrel{{\scriptstyle(ii)}}{{\leq}}\sum_{j=1}^{K}\frac{1}{\lceil Kh_{K}p_{j}/\alpha\rceil}\mathds{1}\left\{h_{K}p_{j}\leq\alpha\right\},$
where $(i)$ is a consequence of $(1/m)\mathds{1}\{Kp_{j}h_{K}\leq\alpha m\}\leq(1/\lceil Kh_{K}p_{j}/\alpha\rceil)\mathds{1}\{Kp_{j}h_{K}\leq\alpha m\}$, for all $j=1,\dots,K$, while $(ii)$ to the fact that $K/m\geq 1$. ∎
Proof of Lemma 7.1.
Let $K\geq 2$ and $T_{K}\geq 1$. Define the function $f:[0,\infty)\to[0,\infty]$ as
$p\mapsto\min\left\{\frac{1}{T_{K}p}-\frac{1}{T_{K}},K\right\}\mathds{1}\{p\in[0,1]\},$
that is decreasing in $p$. Then,
$\begin{split}\int_{0}^{1}f(p)\mathrm{d}p&=\int_{0}^{1}\min\left\{\frac{1}{T_{K}p}-\frac{1}{T_{K}},K\right\}\mathds{1}\{p\in[0,1]\}\mathrm{d}p\\ &=\int_{0}^{(T_{K}K+1)^{-1}}K\mathrm{d}p+\int_{(T_{K}K+1)^{-1}}^{1}\left(\frac{1}{T_{K}p}-\frac{1}{T_{K}}\right)\mathrm{d}p\\ &=\frac{K}{T_{K}K+1}-\frac{K}{T_{K}K+1}+\frac{\log(T_{K}K+1)}{T_{K}}=\frac{\log(KT_{K}+1)}{T_{K}}.\end{split}$
This implies that $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$ if and only if
$KT_{K}+1-e^{T}_{K}\leq 0.$ (37)
We would like to choose $T_{K}$ as small as possible. One possible candidate is $T_{K}=\log K+\log\log K+1$. Indeed, plugging $T_{K}=\log K+\log\log K+1$ into the left-hand side of (37) we find that it is verified when
$\log\log K+1+\frac{1}{K}\leq(e-1)\log K,$
and this holds if $K\geq 2$ by checking $K=2,3$ and using the derivative of both sides for $K\geq 4$. ∎
Proof of Proposition 7.2.
It is straightforward to see that $F$ is an increasing function. By direct calculation,
$\begin{split}\mathbb{P}(F(\mathbf{P})\leq\alpha)&=\mathbb{P}((T_{K}+1)H(\mathbf{P})\leq\alpha)\\ &=\mathbb{P}\left((T_{K}+1)K\left(\sum_{k=1}^{K}\frac{1}{P_{k}}\right)^{-1}\leq\alpha\right)\\ &=\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\alpha}{T_{K}P_{k}}-\frac{1}{T_{K}}\right)\geq 1\right)\\ &\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{T_{K}P_{k}/\alpha}-\frac{1}{T_{K}}\right)\mathds{1}\left\{\frac{P_{k}}{\alpha}\in[0,1]\right\}\geq 1\right)\\ &=\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}\min\left\{\left(\frac{1}{T_{K}P_{k}/\alpha}-\frac{1}{T_{K}}\right),K\right\}\mathds{1}\left\{\frac{P_{k}}{\alpha}\in[0,1]\right\}\geq 1\right)\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\min\left\{\left(\frac{1}{T_{K}P_{k}/\alpha}-\frac{1}{T_{K}}\right),K\right\}\mathds{1}\left\{\frac{P_{k}}{\alpha}\in[0,1]\right\}\right]\\ &=\alpha\mathbb{E}\left[\frac{1}{\alpha}\frac{1}{K}\sum_{k=1}^{K}\min\left\{\left(\frac{1}{T_{K}P_{k}/\alpha}-\frac{1}{T_{K}}\right),K\right\}\mathds{1}\left\{\frac{P_{k}}{\alpha}\in[0,1]\right\}\right]\leq\alpha,\end{split}$
where $(i)$ is due to the fact that $(1/(T_{K}x)-1/T_{K})$ is negative for $x>1$, and $(ii)$ holds due to Markov’s inequality. The last inequality is a consequence of Lemma 2.3 and Lemma 7.1. ∎
Appendix B Exchangeable and randomized p-merging function
In this part, we will integrate the results regarding randomization, introduced in Section 3.3, in the case where the sequence of p-values exhibits exchangeability. In fact, starting from exchangeable p-values, it is possible to prove that if randomization is allowed, then it is possible to improve some results introduced in Section 3.1. We start by defining a “randomized” version of Theorem 3.2.
Theorem B.1.
Let $f$ be a calibrator, and $(\mathbf{P},U)=(P_{1},\dots,P_{K},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$ such that $\mathbf{P}$ is exchangeable. For each $\alpha\in(0,1)$, we have
$\mathbb{P}\left(f\left(\frac{P_{1}}{\alpha}\right)\geq U\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\alpha}\right)\geq 1\right)\leq\alpha.$
Proof.
The proof invokes the exchangeable and uniformly-randomized Markov inequality (EUMI) introduced in Ramdas and Manole, (2023); see Theorem 2.6. In particular,
$\begin{split}\mathbb{P}\left(f\left(\frac{P_{1}}{\alpha}\right)\geq U\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\alpha}\right)\geq 1\right)\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\left[f\left(\frac{P_{1}}{\alpha}\right)\right]=\alpha\mathbb{E}\left[\frac{1}{\alpha}f\left(\frac{P_{1}}{\alpha}\right)\right]\stackrel{{\scriptstyle(ii)}}{{\leq}}\alpha,\end{split}$
where $(i)$ is due to EUMI and $(ii)$ is due Lemma 2.3. ∎
Similarly to how it was done in the preceding sections, let us now define an exchangeable and randomized p-merging function.
Definition B.2.
An exchangeable and randomized p-merging function is an increasing Borel function $F:[0,1]^{K+1}\to[0,1]$ such that $\mathbb{P}(F(\mathbf{P},U)\leq\alpha)\leq\alpha$ for all $\alpha\in(0,1)$ and $(\mathbf{P},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$ with $\mathbf{P}$ exchangeable. It is homogeneous if $F(\gamma\mathbf{p},u)=\gamma F(\mathbf{p},u)$ for all $\gamma\in(0,1]$ and $(\mathbf{p},u)\in[0,1]^{K+1}$. An exchangeable and randomized p-merging function is admissible if for any exchangeable and randomized p-merging function $G$, $G\leq F$ implies $G=F$.
Let $f$ be a calibrator; then for $\alpha\in(0,1)$, we define the exchangeable and randomized rejection region by
$R_{\alpha}=\left\{(\mathbf{p},u)\in[0,1]^{K+1}:f\left(\frac{p_{1}}{\alpha}\right)\geq u\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\right\},$
where we set $f(p_{i}/u)=0$ if $u=0$. Starting from $R_{\alpha}$, we can define the function $F:[0,1]^{K+1}\to[0,1]$ by
$\begin{split}F(\mathbf{p},u)&=\inf\left\{\alpha\in(0,1):(\mathbf{p},u)\in R_{\alpha}\right\}\\ &=\inf\left\{\alpha\in(0,1):f\left(\frac{p_{1}}{\alpha}\right)\geq u\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\geq 1\right\},\end{split}$ (38)
with the convention $\inf\varnothing=1$ and $0\times\infty=\infty$.
Theorem B.3.
If $f$ is a calibrator and $(\mathbf{P},U)\in\mathcal{U}^{K}\otimes\mathcal{U}$ with $\mathbf{P}$ exchangeable, then $F$ in (38) is a homogeneous exchangeable and randomized p-merging function.
Proof.
It is clear that $F$ is increasing and Borel since $R_{\alpha}$ is a lower set. For an exchangeable $\mathbf{P}\in\mathcal{U}^{K}$ and $\alpha\in(0,1)$, using Theorem B.1 and the fact that $(R_{\beta})_{\beta}\in(0,1)$ is nested, we have
$\begin{split}\mathbb{P}(F(\mathbf{P},U)\leq\alpha)&=\mathbb{P}\left(\inf\left\{\beta\in(0,1):f\left(\frac{P_{1}}{\beta}\right)\geq U\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\geq 1\right\}\leq\alpha\right)\\ &=\mathbb{P}\left(\bigcap_{\beta>\alpha}\left\{f\left(\frac{P_{1}}{\beta}\right)\geq U\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\geq 1\right\}\right)\\ &=\inf_{\beta>\alpha}\mathbb{P}\left(f\left(\frac{P_{1}}{\beta}\right)\geq U\mathrm{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\exists k\leq K:\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{P_{i}}{\beta}\right)\geq 1\right)\\ &\leq\inf_{\beta>\alpha}\beta=\alpha.\end{split}$
therefore $F$ is a valid exchangeable and randomized p-merging function. Homogeneity comes directly from the definition of (38). ∎
Appendix C Generalized Hommel combination rule
We can generalize the Hommel combination by allowing the selection of certain quantiles from the possible $K$ different quantiles of $\mathbf{p}=(p_{1},\dots,p_{K})$, for example, one can select the minimum between $K$ times the minimum, 2 times the median and the maximum. This can be obtained by noting that the Hommel combination rule in (16) can be rewritten in terms of quantiles. In particular, the following holds:
$F^{\prime}_{\mathrm{Hom}}(\mathbf{p}):=h_{K}\bigwedge_{k=1}^{K}\frac{1}{\lambda_{k}}p_{(\lceil\lambda_{k}K\rceil)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{with }h_{K}=\sum_{j=1}^{K}\frac{\lambda_{j}-\lambda_{j-1}}{\lambda_{j}},$
where $(\lambda_{0},\lambda_{1},\dots,\lambda_{K})$ is the vector of quantiles such that $\lambda_{j}=j/K,\,j=0,1,\dots,K$. This gives the intuition to define a generalization of the aforementioned rule.
Let us define the vector of quantiles $\lambda=(\lambda_{0},\lambda_{1},\dots,\lambda_{M})$, such that $\lambda_{0}=0,\,\lambda_{j}\in(0,1],$ if $j=1,\dots,M$ and $\lambda_{j}<\lambda_{j+1}$. Then, we can define
$F^{\prime}_{\mathrm{GHom}}(\mathbf{p}):=h_{M}\bigwedge_{k=1}^{M}\frac{1}{\lambda_{k}}p_{(\lceil\lambda_{k}K\rceil)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{with }h_{M}=\sum_{j=1}^{M}\frac{\lambda_{j}-\lambda_{j-1}}{\lambda_{j}}.$ (39)
Lemma C.1.
Let $f$ be a function defined by
$f(p)=\sum_{j=1}^{M}\frac{1}{\lambda_{j}}\mathds{1}\left\{p\in\left(\frac{\lambda_{j-1}}{h_{M}},\frac{\lambda_{j}}{h_{M}}\right]\right\}+\infty\mathds{1}\{p=0\}.$ (40)
Then $f$ is an admissible calibrator. Moreover, the p-merging function induced by $f$ is
$F_{\mathrm{GHom}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{M}\frac{1}{\lambda_{j}}\mathds{1}\left\{\frac{p_{k}}{\alpha}\in\left(\frac{\lambda_{j-1}}{h_{M}},\frac{\lambda_{j}}{h_{M}}\right]\right\}\geq 1\right\},$ (41)
is valid and it dominates (39).
Proof.
It is simple to see that $f$ is decreasing and upper semicontinuous. In addition,
$\int_{0}^{1}f(p)\mathrm{d}p=\sum_{j=1}^{M}\frac{1}{\lambda_{j}}\left(\frac{\lambda_{j}-\lambda_{j-1}}{h_{M}}\right)=\frac{1}{h_{M}}\sum_{j=1}^{M}\frac{\lambda_{j}-\lambda_{j-1}}{\lambda_{j}}=1.$
This implies that $F_{\mathrm{GHom}}$ is admissible. To prove the last part, we see that for any $\mathbf{p}\in(0,1]^{K}$ and $\alpha\in(0,1)$ we have that $F_{\mathrm{GHom}}(\mathbf{p})\leq\alpha$, if and only if,
$\bigwedge_{k=1}^{M}\frac{1}{\lambda_{k}}p_{(\lceil\lambda_{k}K\rceil)}\leq\frac{\alpha}{h_{M}}.$
This implies that, for some $m\in\{1,\dots,K\}$, we have that
$\sum_{i=1}^{K}\frac{1}{K}\mathds{1}\left\{p_{i}\leq\alpha\frac{\lambda_{m}}{h_{M}}\right\}\geq\lambda_{m}.$
We now define this chain of inequalities,
$\begin{split}1&\leq\sum_{i=1}^{K}\frac{1}{K\lambda_{m}}\mathds{1}\left\{p_{i}\leq\alpha\frac{\lambda_{m}}{h_{M}}\right\}=\frac{1}{K}\sum_{i=1}^{K}\frac{1}{\lambda_{m}}\sum_{j=1}^{m}\mathds{1}\left\{\frac{p_{i}}{\alpha}\in\left(\frac{\lambda_{j-1}}{h_{M}},\frac{\lambda_{j}}{h_{M}}\right]\right\}\\ &\leq\frac{1}{K}\sum_{i=1}^{K}\sum_{j=1}^{m}\frac{1}{\lambda_{j}}\mathds{1}\left\{\frac{p_{i}}{\alpha}\in\left(\frac{\lambda_{j-1}}{h_{M}},\frac{\lambda_{j}}{h_{M}}\right]\right\}\leq\frac{1}{K}\sum_{i=1}^{K}\sum_{j=1}^{M}\frac{1}{\lambda_{j}}\mathds{1}\left\{\frac{p_{i}}{\alpha}\in\left(\frac{\lambda_{j-1}}{h_{M}},\frac{\lambda_{j}}{h_{M}}\right]\right\}.\end{split}$
It is possible to see that (39) is a special case of the combination defined in Rüger, (1978) when $M=1$, while it coincides with the Hommel combination rule when $\lambda=(0,1/K,\dots,(K-1)/K,1)$.
Appendix D Improving generalized mean
In this section, we discuss the generalized mean combination rule defined in (13), for $r\in\mathbb{R}\setminus\{0\}$. This combination rule, introduced in Vovk and Wang, (2020), is quite broad and contains some important cases well known in the literature. In particular, if $r=1$ it reduces to the sample average (Section 6), while if $r=-1$ it coincides with the harmonic mean described in Section 7. We introduce a lemma characterizing the calibrator used in the context of the generalized mean combination rule.
Lemma D.1.
Let $r\in\mathbb{R}\setminus\{0\}$ and $f:[0,\infty)\to[0,\infty]$ be given by
$f(p)=\min\left\{\frac{r(1-p^{r})}{T_{r,K}},K\right\}\mathds{1}\{p\in[0,1]\},$ (42)
where $T_{r,K}>0$ is any constant, possibly dependent on $K$, such that $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$. Then $f$ is a calibrator.
Proof.
Since $f(p)$ is continuously decreasing in $T_{r,K}$, it can be verified that there exists $T_{r,K}>0$ such that $\int_{0}^{1}f(p)\mathrm{d}p=1$. Moreover, $f$ is decreasing and non-negative which completes the claim. ∎
It is simple to see that for $r>0$, we have that $T_{r,K}=r^{2}/(r+1)$ satisfies $\int_{0}^{1}f(p)\mathrm{d}p$. From previous sections, we obtain $T_{1,K}=1/2$ while a more complex result appears for $T_{-1,K}$. We now see how the calibrator defined in (42) is related to the rule defined in (13). First, we define the function $F^{\prime}_{\mathrm{M_{r}}}$ as
$F^{\prime}_{\mathrm{M_{r}}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\frac{r(1-(p_{k}/\alpha)^{r})}{T_{r,K}}\geq 1\right\}=\frac{M_{\mathrm{r}}(\mathbf{p})}{(1-T_{r,K}/r)^{1/r}},$ (43)
where $M_{\textrm{r}}(\mathbf{p})$ is the $r$-generalized mean of $\mathbf{p}$, defined by $M_{\textrm{r}}(\mathbf{p})=(\sum_{k=1}^{K}p_{k}^{r}/K)^{1/r}$. In particular, $F^{\prime}_{\mathrm{M_{r}}}(\mathbf{p})$ coincides with the rule defined (13) where $a_{r,K}=(1-T_{r,K}/r)^{-1/r}$. In addition, if $r>0$ then $F^{\prime}_{\mathrm{M_{r}}}(\mathbf{p})=(r+1)^{1/r}{M_{r}}(\mathbf{p})$ which coincides with the asymptotically precise merging function studied by Vovk and Wang, (2020). It is possible to prove that $F^{\prime}_{\mathrm{M_{r}}}(\mathbf{p})\leq F_{\mathrm{M_{r}}}(\mathbf{p})$, where
$F_{\mathrm{M_{r}}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(\frac{r(1-(p_{k}/\alpha)^{r})}{T_{r,K}}\right)_{+}\geq 1\right\},$ (44)
which is the p-merging function induced by the calibrator defined in (42). In particular, according to Theorem 2.2 we have that $F_{\mathrm{M_{r}}}(\mathbf{p})$ is a valid p-merging function.
D.1 Exchangeable generalized mean
We now define the function $F_{\mathrm{EM_{r}}}$ in the following way:
$F_{\mathrm{EM_{r}}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(\frac{r(1-(p_{i}/\alpha)^{r})}{T_{r,K}}\right)_{+}\geq 1\right\}.$ (45)
If we define $F^{\prime}_{\mathrm{EM_{r}}}$ by
$F^{\prime}_{\mathrm{EM_{r}}}(\mathbf{p})=\inf\left\{\alpha\in(0,1):\exists\ell\leq K\mathrm{\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ }\frac{1}{\ell}\sum_{i=1}^{\ell}\left(\frac{r(1-(p_{i}/\alpha)^{r})}{T_{r,K}}\right)\geq 1\right\},$ (46)
then it holds that $F_{\mathrm{EM_{r}}}\leq F^{\prime}_{\mathrm{EM_{r}}}$.
Theorem D.2.
Let $r\in\mathbb{R}\setminus\{0\}$, then the function $F^{\prime}_{\mathrm{EM_{r}}}$ defined in (46) equals
$\left(1-\frac{T_{r,K}}{r}\right)^{-1/r}\left(\bigwedge_{m=1}^{K}M_{r}(\mathbf{p}_{m})\right),$
is a valid exchangeable p-merging function, and it strictly dominates the function $F_{\mathrm{M_{r}}}$ in (44). However, it is strictly dominated by the function $F_{\mathrm{EM_{r}}}$ defined in (45) that is also a valid exchangeable p-merging function.
Proof.
According to Theorem 3.4 and Lemma D.1, it follows that the function $F_{\mathrm{EM_{r}}}$ is a valid exchangeable p-merging function. In addition, fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$, then $F^{\prime}_{\mathrm{EM_{r}}}\leq\alpha$ if and only if
$\frac{1}{\ell}\sum_{i=1}^{\ell}\frac{r(1-(p_{i}/\alpha)^{r})}{T_{r,K}}\geq 1\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }\ell\leq K\implies\left(1-\frac{T_{r,K}}{r}\right)^{-1/r}\left(\frac{1}{\ell}\sum_{i=1}^{\ell}p_{i}^{r}\right)^{1/r}\leq\alpha\mathrm{\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ }\ell\leq K.$
Taking a minimum over $\ell$ yields the desired formula. ∎
D.2 Randomized generalized mean
According to the the previous sections, we define the randomized p-merging function $F_{\mathrm{UM_{r}}}$ as follows
$F_{\mathrm{UM_{r}}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(\frac{r(1-(p_{k}/\alpha)^{r})}{T_{r,K}}\right)_{+}\geq u\right\}.$ (47)
The function $F_{\mathrm{UM_{r}}}\leq F^{\prime}_{\mathrm{UM_{r}}}$ where the function $F^{\prime}_{\mathrm{UM_{r}}}$ is defined by
$F^{\prime}_{\mathrm{UM_{r}}}(\mathbf{p},u)=\inf\left\{\alpha\in(0,1):\frac{1}{K}\sum_{k=1}^{K}\left(\frac{r(1-(p_{k}/\alpha)^{r})}{T_{r,K}}\right)\geq u\right\}=\frac{M_{r}(\mathbf{p})}{(1-uT_{r,K}/r)^{1/r}},$ (48)
and can be considered as the randomized version of (43).
Theorem D.3.
Let $r\in\mathbb{R}\setminus\{0\}$, then the function $F^{\prime}_{\mathrm{EM_{r}}}$ defined in (48) equals
$\frac{M_{r}(\mathbf{p})}{(1-uT_{r,K}/r)^{1/r}},$
is a valid exchangeable p-merging function, and it strictly dominates the function $F_{\mathrm{M_{r}}}$ in (44). However, it is strictly dominated by the function $F_{\mathrm{UM_{r}}}$ defined in (47) that is also a valid exchangeable p-merging function.
Proof.
According to Corollary 3.13 and Lemma D.1, it follows that $F_{\mathrm{UM_{r}}}$ is a valid randomized p-merging function. In addition, fix any $\alpha\in(0,1)$ and $\mathbf{p}\in(0,1]^{K}$, then $F^{\prime}_{\mathrm{UM_{r}}}\leq\alpha$ if and only if
$\frac{1}{K}\sum_{k=1}^{K}\left(\frac{r(1-(p_{k}/\alpha)^{r})}{T_{r,K}}\right)\geq u\implies\frac{M_{r}(\mathbf{p})}{(1-uT_{r,K}/r)^{1/r}}\leq\alpha.$
Appendix E General algorithm
One general algorithm to compute the exchangeable p-merging defined in (6) induced by a calibrator $f$ is proposed in Algorithm 1. The algorithm employs the bisection method and it consistently yields a p-value exceeding that of the induced exchangeable p-merging function by at most $2^{-B}$.
Data: A calibrator $f$, $B\in\mathbb{N}$, and a sequence of p-values $(p_{1},\dots,p_{K}).$
$L:=0$ and $U:=1$;
for $i=1,\dots,B$ do
       $\alpha:=(L+U)/2$;
       if $\bigvee_{k=1}^{K}\left(\frac{1}{k}\sum_{i=1}^{k}f\left(\frac{p_{i}}{\alpha}\right)\right)\geq 1$ then
             $U:=\alpha$
       end if
      else
             $L:=\alpha$
       end if
end for
Result: $U$
Algorithm 1 Exchangeable p-merging function
Appendix F Additional simulation results
In this section we report some additional simulation results. We first compare the 3 different combination rules defined in Section 6.2, respectively, $F_{\mathrm{UA}},F^{\prime}_{\mathrm{UA}},F^{*}_{\mathrm{UA}}$. In particular, we recall that $F_{\mathrm{UA}}$ defined in (25) dominates $F^{\prime}_{\mathrm{{UA}}}$ defined in (26), while the combination $F^{*}_{\mathrm{UA}}$ (introduced in Wang, (2024, Appendix B.2)) neither dominates nor is dominated by either of the two.
Figure 3: Combination of p-values using different randomized combination rules based on the average of p-values (and the Bonferroni method, for comparison). $F^{*}_{\mathrm{UA}}$ is more powerful than $F^{\prime}_{\mathrm{UA}}$ only when $\mu\lesssim 2.5$.
The set of $K=100$ p-values is generated as in (36), and two different values of $\rho$ are chosen, 0.9 and 0.1, respectively. The results are obtained by repeating the procedure $10,000$ times and reporting the average. From Figure 3, we can see that the power $F_{\mathrm{UA}}$ is always higher than the power of the other two combination rules. The function $F^{*}_{\mathrm{UA}}$ is more powerful than $F^{\prime}_{\mathrm{UA}}$ only in the first part of the $y$-axis in both cases (for $\mu\leq 2.5$, indicatively), so there is no clear preference between the two. As expected, the Bonferroni correction is more powerful near independence ($\rho=0.1$).
In addition, we compare all the randomized combination rule reported in the paper. We omit the randomized functions that are dominated by other randomized p-merging functions. In this case, $K=100$ p-values are generated as in (36), and two different values of $\rho$ are chosen, 0.9 and 0.1, respectively. The parameter $k$ for the randomized Rüger combination rule is set to $K/2$ and $\mu$ varies in the interval $[0,3]$. The results are obtained by repeating the procedure $10,000$ times and reporting the average.
Figure 4: Combination of p-values using different randomized p-merging functions. The order of the performance of the different exchangeable p-merging functions is almost the opposite in the two situations.
The results of the randomized versions of “twice” the median and (improved) “twice” the mean are similar in both scenario. In addition, $F_{\mathrm{UHom}}$, defined in (20), and $F_{\mathrm{UH}}$, defined in (29) are quite similar in terms of power (Figure 4). From Figure 4, we can observe an opposite behavior of the functions in the case where $\rho=0.9$ or $\rho=0.1$. In general, the most powerful functions in the left graph tend to be the least powerful in the right graph, and vice versa.